### v1

# From Evaluation to Enhancement: Large Language Models for Zero-Knowledge Proof Code Generation

Abstract.
Zero-knowledge proofs (ZKPs) are increasingly deployed in domains such as privacy-preserving authentication, blockchain scalability, and secure finance. However, authoring ZK programs remains challenging: unlike mainstream programming, ZK development requires reasoning about finite field arithmetic, constraint systems, and gadgets, making it knowledge-intensive and error-prone. While large language models (LLMs) have demonstrated strong code generation capabilities in general-purpose languages, their effectiveness for ZK programming, where correctness hinges on both language mastery and gadget-level reasoning, remains unexplored. To address this gap, we propose ZK-Eval, a domain-specific evaluation pipeline that probes LLM capabilities at three levels: language knowledge, gadget competence, and end-to-end program generation. Our evaluation of four state-of-the-art LLMs reveals that models excel at surface-level syntax but struggle with gadget usage and semantic correctness, often yielding incorrect programs. Based on these insights, we introduce ZK-Coder, an agentic framework that augments LLMs with constraint sketching, guided retrieval, and interactive repair. Experiments on Circom and Noir show substantial gains, with success rates improving from 17.35% to 83.38% and from 32.21% to 90.05%, respectively. With ZK-Eval and ZK-Coder, we establish a foundation for systematically measuring and augmenting LLMs in ZK code generation to lower barriers for practitioners and advance trustworthy computation.
Zero-Knowledge Proofs, Large Language Models, Code Generation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xjournalyear: 2025publicationmonth: 9copyright: noneccs: Software and its engineering Automatic programmingccs: Security and privacy Cryptographyccs: Software and its engineering Domain specific languagesccs: Computing methodologies Natural language processing
1. Introduction
Zero-knowledge proofs (ZKPs) are a powerful cryptographic primitive that allow one party to prove knowledge of a statement without revealing the underlying witness. In practice, this means proving that a computation or claim is correct while keeping sensitive inputs private. ZKPs have moved far beyond theory: they are now central to privacy-preserving authentication, blockchain scalability, and emerging applications across finance and security domains (Gupta, 2025; Lavin et al., 2024; team, 2024). As demand for trustworthy computation grows, the ability to author, test, and verify ZKP programs has become a first-class concern in the software engineering community (Takahashi et al., 2025; Hochrainer et al., 2024; Xu et al., 2025; Pailoor et al., 2023; Jiang et al., 2025a).
Figure 1. Illustrative Comparison Between the Process of Mainstream and ZK Programming.
However, authoring ZK programs is challenging. Unlike mainstream programming, which builds on the von Neumann model and imperative execution, ZKP programs boil down to specifying and proving mathematical constraints over finite fields. Its development requires reasoning about finite field arithmetic, constraint systems, and gadgets — a novel concept advocated by modern ZKP frameworks to encode reusable, domain-specific building blocks. As shown in Fig. 1 and Table 1, ZK programs do not describe how to execute a computation but how to prove its correctness by compiling constraints into arithmetic circuits. Developers must specify equations over inputs, outputs, and intermediate variables, carefully wiring gadgets to enforce correctness. Even small omissions or miswirings can yield satisfiable but unsound constraint systems, leading to errors that compilers rarely flag. This makes ZK development knowledge-intensive, error-prone, and inaccessible to non-experts, despite significant progress in proof systems and verification tools (Chaliasos et al., 2024; Pailoor et al., 2023; Takahashi et al., 2025).
Table 1. Comparison of Mainstream and ZK Programming.
Aspect Mainstream ZK Program
Paradigm Imperative/OOP/Func Constraint Systems
Comp. Model von Neumann Arch Arithmetic Circuits
Purpose to Compute to Prove
Data Types Int/Float/Str/… Finite Field Elements
Mutability Dynamic Static
Turing-complete? Yes No
Control Flow If/Loop/Recursion Limited
Side Effects Yes No, Pure Math
Figure 2. Overview of ZK-Eval and ZK-Coder.
As a result, ZK programming demands significantly more time and expertise than mainstream software development, posing a steep barrier to entry for many practitioners. In contrast, recent advances in large language models (LLMs) have lowered this barrier in general-purpose languages by demonstrating strong code generation capabilities (Jiang et al., 2025b), achieving impressive results on benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021), and powering widely adopted tools like GitHub Copilot (GitHub, 2025) and Cursor (team, 2025). This contrast naturally raises the question: to what extent can modern LLMs assist in generating ZK code from natural language specifications? Answering this question requires more than surface-level benchmarking. Compared to programming in mainstream languages, success in ZK development hinges on two competencies: (i) knowledge of domain-specific languages (DSLs) and their toolchains, and (ii) mastery of gadgets as building blocks for constraint systems. Existing benchmarks focus only on mainstream languages and overlook these unique requirements.
To address this gap, we introduce ZK-Eval, a domain-specific evaluation pipeline that probes LLM capability at three levels: language knowledge, gadget-level competence, and end-to-end program code generation (left panel of Fig. 2). We anticipate that this design would provide a structured way to pinpoint where models succeed, where they fail, and why, rather than treating code generation as a monolithic task. Applying ZK-Eval to four state-of-the-art LLMs (including advanced reasoning models like GPT-o4-mini and GPT-o3) reveals a clear pattern. While models perform strongly on surface-level language knowledge, their accuracy drops significantly when reasoning about gadgets or assembling complete programs, often producing code that compiles but fails semantically. These results underscore both the promise of LLMs in lowering barriers to ZK development and the need for augmentations to bridge the gap between syntactic knowledge and reliable code generation. Motivated by these findings, we present ZK-Coder (right panel of Fig. 2), an agentic framework that augments LLMs with constraint sketching, retrieval, and interactive repair. By grounding generation in constraint reasoning and guided gadget usage, ZK-Coder achieves substantial gains in reliability, improving success rates from 17.35% to 83.38% on Circom and from 32.21% to 90.05% on Noir. In summary, our contributions are as follows:
  1. (1)
    We propose ZK-Eval, a three-stage evaluation pipeline for systematically assessing LLM capabilities in ZK code generation.
  2. (2)
    We introduce ZK-Coder, an agentic framework that augments LLMs with constraint sketching, guided retrieval, and interactive repair to improve reliability.
  3. (3)
    We evaluate the performance of four state-of-the-art LLMs on ZK-Eval and demonstrate how ZK-Coder substantially improves end-to-end code generation.
2. Background
2.1. Zero-Knowledge Proofs and Programming Frameworks
Zero-knowledge proofs (ZKPs) are cryptographic protocols that allow one party to prove to another that a statement is true without revealing any information beyond the validity of the statement itself. Formally, the following properties are desired for a correct ZKP program.
  1. (1)
    Completeness: If a statement is true, an honest prover can convince the verifier of its validity.
  2. (2)
    Soundness: If a statement is false, no dishonest prover can convince the verifier of its validity.
  3. (3)
    Zero-Knowledge: If a statement is true, the verifier learns nothing beyond the validity of the statement itself.
These properties make ZKPs particularly useful in privacy-preserving applications (Gupta, 2025; team, 2024; Lavin et al., 2024). However, achieving these properties relies on the correct implementation of the underlying cryptographic primitives and protocols with respect to the high-level intent. Consequently, writing ZKP programs can be challenging due to the complexity of the underlying field-arithmetic mathematics and the need for specialized knowledge in cryptography (Pailoor et al., 2023; Wen et al., 2024b; Chaliasos et al., 2024).
Table 2. Tools for Programming Zero-Knowledge Proofs
Frontend DSL Compiler GitHub repository
Circom (Bellés-Muñoz et al., 2022) Circom Rust iden3/circom (1.5K)
Noir (Aztec Labs, 2025) Rust-like Rust noir-lang/noir (1.2K)
Leo (Chin et al., 2021) Rust-like Rust AleoHQ/leo (4.8K)
ZoKrates (Contributors, 2025c) Python-like Rust Zokrates/ZoKrates (1.9K)
Loosely speaking, developers need to translate their high-level intent into an arithmetic representation $R(x,w)=1$, where $x$ denotes public inputs, $w$ denotes the private witness, and $R$ is a collection of polynomial equations over field elements. To ease the development process, several programming frameworks have emerged, providing higher-level abstractions and DSLs for expressing ZKP computations (Sheybani et al., 2025; Ozdemir et al., 2022; Thomas, 2022). In Table 2, we review several representative frontend frameworks, including Circom, Noir, and others. These frameworks differ not only in syntax but also in how they expose the underlying constraint system, which directly impacts the difficulty of automatic code generation. Circom requires developers to think in terms of low-level circuit wiring and explicit constraints, which poses challenges for LLM-based generation since the model must produce correct and efficient circuit templates (Hochrainer et al., 2024; Chaliasos et al., 2025; Takahashi et al., 2025). Noir, by contrast, adopts a Rust-like syntax and enforces type safety, making it closer to conventional programming languages and thus more approachable for LLMs trained on general-purpose code corpora (Joel et al., 2024; Jiang et al., 2024b, 2025b). As a concurrent effort, there are also emerging frameworks that target general-purpose languages, such as SP1 (Labs, 2025), Nexus (Contributors, 2025a), and RISC Zero (Contributors, 2025b), which compiles Rust code into a restricted instruction set and generates ZK proofs over the execution trace within a virtual machine. However, these frameworks often come with significant performance overheads due to the complexity of instruction emulation within the virtual machine. Therefore, we consider them as an orthogonal direction to our work for reducing the burden of writing ZKP programs. Taken together, these tools illustrate the diverse design space of ZKP programming frontends, and highlight why aligning LLM-based code generation with specific framework characteristics is essential for usability and correctness.
In this work, we focus on Circom and Noir for two main reasons. First, they are the most widely adopted general-purpose ZK frontends in both academia and industry, with active communities and production deployments (e.g., Circom in Ethereum-based rollups, Noir in Aztec). This ensures that advances in LLM-assisted code generation for these languages can have immediate practical impact. Second, they represent two complementary design paradigms that are highly relevant for evaluating code generation: Circom exposes low-level, circuit-oriented constructs where the challenge for LLMs is to correctly translate high-level intent into constraint wiring, whereas Noir provides a Rust-like syntax with strong typing and higher-level abstractions, which better aligns with the training data of current LLMs. By studying both, we capture the spectrum from circuit-level specification to program-level ergonomics, allowing us to analyze how LLMs handle different abstraction levels in ZKP development.
2.2. LLM-based Code Generation
Large Language Models (LLMs) have emerged as powerful assistants in software development, enabling code generation from natural language prompts with tools like GitHub Copilot, Cursor, and model-driven agents quickly translating intent into code (Jiang et al., 2024b, 2025b). These systems are reshaping engineering workflows, aiding in writing, testing, and documenting code by drawing on vast amounts of publicly available repositories and documentation (GitHub, 2025).
Framework.
Modern LLM-based code generation typically follows an agentic framework (Zhang et al., 2024a; Wang et al., 2025), where the model acts as an intelligent assistant, interpreting user intent, retrieving relevant information, and producing code snippets or entire functions as needed. This process often involves multiple iterations, with the model refining its output based on user feedback and additional context. In this way, it can leverage off-the-shelf LLMs effectively without code-specific fine-tuning (Dong et al., 2025; Pan et al., 2025). However, to the best of our knowledge, there has been limited exploration of LLMs for code generation in the context of low-resource or DSLs such as those used in ZKPs.
Benchmark.
As an important part of LLM-based code generation research, benchmarks play a crucial role in understanding model performance and guiding future improvements. Currently, their accuracy and reasoning ability are usually evaluated on standard benchmarks on mainstream programming languages such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021). However, evaluation tools for low-resource and DSLs remain limited, hindering comprehensive assessments of LLM performance in these contexts (Joel et al., 2024; Giagnorio et al., 2025). Code generation for ZKP programs exemplifies the challenges and opportunities in this space. Though benchmarks for low-resource languages (Cassano et al., 2022; Yang et al., 2023) and DSLs (Liu et al., 2023; Abukhalaf et al., 2023) exist, to our knowledge, no standardized benchmark suite specifically targets ZKP code generation, requiring tailored benchmarks and evaluation strategies.
3. Measuring LLM Capability for ZK Code Generation
As discussed in Sec. 1 and Sec. 2.1, ZK programming follows a proof-oriented model: developers translate high-level specifications into algebraic constraints over a finite field and produce proofs of satisfiability. This shift yields the workflow depicted in Fig. 1, from constraint specification to implementation by composing gadgets and finally proof generation. We therefore assess two capabilities essential for end-to-end success: ➀ language and toolchain knowledge and ➁ gadget-level competence.
  1. (1)
    Language and Toolchain Knowledge. Developers must understand syntax, typing rules, semantics of field operations, and the structure of witness assignments. Weakness in any of these areas routinely leads to programs that cannot be compiled, or programs that compile but encode an invalid relation.
  2. (2)
    Gadget-level Competence. Many basic operations in ZKP programs are not natively attainable in the programming frameworks. Instead, developers must rely on a library of pre-defined gadgets that encapsulate these operations with optimized implementations. Gadget-level competence refers to the ability to effectively utilize these gadgets in the context of ZKP development and, when necessary, implement new ones from scratch.
Taken together, these two capabilities form the foundation of end-to-end generation of ZK programs. Without reliable language and toolchain knowledge, LLMs are unable to produce compilable programs, while without gadget-level competence, the resulting programs cannot faithfully encode the intended relations. In practice, developers move between applying syntax and compilation rules to structure programs and assembling or extending gadgets to realize non-trivial constraints, and for LLMs, success also hinges on integrating these skills that bridge surface-level language fluency with the deeper ability to model arithmetic and logical constraints. Existing benchmarks, however, primarily target mainstream programming models, leaving this special integration of capabilities unexplored. This gap motivates our study: to systematically measure and evaluate LLM capability for ZK code generation. To study that, our evaluation pipeline introduces ZK-Eval, which explicitly measures both dimensions and assesses how they combine in end-to-end generation tasks.
3.1. ZK-Eval Benchmark
In this section, we present the curation of ZK-Eval. Instead of focusing solely on end-to-end program generation, we aim to measure the individual competencies that contribute to successful ZK code generation. Below, we outline the key components of ZK-Eval.
Figure 3. The Design of MCQ Benchmark Evaluating LLM’s Knowledge on ZK Languages.
3.1.1. Language and Toolchain Knowledge
We first evaluate whether LLMs possess the foundational knowledge required to author programs in ZK languages. To construct this evaluation, we curate a knowledge base covering the two most widely used languages, Circom and Noir. As shown in the left side of Fig. 3, sources are collected through automated web crawling of official documentation, including language specifications, tutorials, and API references (e.g., circomlib (iden3, 2025) for Circom and the Noir standard library). In addition, we manually gather publicly available examples from GitHub repositories (TEZCAN, 2025b; Sirin, [n. d.]; Contributors, [n. d.]; Thor, [n. d.]) and online tutorials (TEZCAN, 2025a) to ensure coverage of idiomatic usage.
Based on this knowledge base, as shown in the middle of Fig. 3, we design a pipeline to automatically generate multiple-choice questions, a format widely adopted in prior evaluations of language models (Hendrycks et al., 2020; Myrzakhan et al., 2024; Zellers et al., 2019; Clark et al., 2018). Inspired by earlier work on LLM-driven question generation (Liu et al., 2020; Abbasiantaeb et al., 2024; Long et al., 2024; Ji et al., 2025), our pipeline iteratively samples subsets of documents and provides them to a large model, which composes candidate questions and answers. This process is repeated until the full document collection has been exhaustively covered, producing a diverse set of drafts.
All generated questions are then subjected to rigorous expert validation (right side of Fig. 3). Human reviewers examine each question for factual correctness, clarity, and alignment with the intended level of difficulty. Cross-checking is performed among the authors to ensure agreement, and questions that do not meet these standards are either corrected or discarded. Topic coverage is also verified to ensure that all major aspects of the knowledge base are represented. We categorize the questions into four groups: basic syntax (language grammar such as datatypes, statements, operators, control flows, signals and buses), advanced topics (e.g., taggings, traits, generics), API references (standard library templates and functions), and compiler principles (translation from source code to arithmetic circuits). A summary of the questions and representative examples is shown in Fig. 4. As shown, the final benchmark consists of 172 questions for Circom and 164 for Noir, with a balanced distribution across the four categories. The full set of questions is provided in the Artifact.
Figure 4. Illustrative MCQ on Basic Circom Syntax with a Summary of Knowledge Categories.
Figure 5. The Design of Gadget Benchmark on LLM’s Ability to Use or Implement Gadgets to Encode Constraints.
3.1.2. Using and Implementing Gadgets to Encode Constraints
The second dimension of ZK-Eval assesses the ability of LLMs to use and implement gadgets to encode constraints, which are the fundamental building blocks of ZK programs. To construct this evaluation, we curate a representative set of gadgets from three complementary sources, as shown in Fig. 5. First, circomlib (iden3, 2025), the official standard library of Circom, provides canonical implementations of common gadgets that developers routinely instantiate. Second, zk-kit (Nico Serrano, [n. d.]) extends this idea to a broader ecosystem, offering more reusable gadget implementations across multiple ZK languages. Third, as an important observation, many gadgets share similar logical and arithmetic structures with primitives in SMT solvers. To capture this perspective, we leverage the Z3Py (Research, [n. d.]) interface to the Z3 SMT solver (De Moura and Bjørner, 2008). These sources span both language-specific and language-agnostic perspectives, ensuring that our benchmark captures the core constraint patterns encountered in practice.
The resulting gadget set consists of 35 gadgets covering a wide range of categories: (i) basic checks, including range tests and comparisons; (ii) logical gates, such as AND, OR, XOR, and equality; (iii) arithmetic operators, including addition, subtraction, multiplication, division, exponentiation, modular reduction, and inversion; and (iv) composite operators, such as sum, distinct, and multiplexors (see the full list in the Supplementary Material). This collection is sufficiently expressive to simulate the primitive layer of typical ZK development workflows: any non-trivial circuit can be realized through the composition of these gadgets.
For every gadget in the set, we author a concise task description. Test cases are crafted to validate both completeness: accepting input assignments and soundness: rejecting assignments. For logical gates, this corresponds to enumerating the truth table; for arithmetic and composite gadgets, it is manually constructed based on their semantics. These tests ensure that correctness is evaluated not only in terms of producing the right outputs, but also in enforcing the intended constraints.
Figure 6. The Design of End-to-end ZK Program Generation Benchmark Adapted from HumanEval.
3.1.3. End-to-End ZK Program Generation
The third dimension of ZK-Eval evaluates the capability of LLMs to perform full end-to-end ZK program generation, which requires the model to integrate language proficiency and gadget-level competence to assemble complete programs that encode non-trivial problem specifications. We build the benchmark on the top of the widely used HumanEval dataset (Chen et al., 2021) with ZKP-specific adaptions, as shown in Fig. 6. In particular, we exclude tasks involving (i) string manipulation, (ii) floating-point arithmetic, and (iii) outputs of variable length. Strings and floating-point numbers are not natively supported by Circom or Noir, while variable-length outputs are incompatible with the static nature of arithmetic circuits. Following prior work (Myrzakhan et al., 2024), each task is automatically annotated by an LLM and verified by experts to ensure accurate filtering, yielding a set of 68 standalone tasks suitable for ZK program generation.
To reflect the verification nature of ZK programs, we further reformulate the problem statements. Original HumanEval tasks are expressed as construction problems, in which a function computes an output $S=P(I)$ for input $I$. In contrast, ZK programs implement verification problems, where the goal is to check whether a candidate solution $S$ satisfies the intended relation for $I$. We therefore restate each task as a verification function $V(I,S)\in\{0,1\}$, which outputs one if and only if the candidate solution is correct.
For each task, we construct an executable Python oracle that implements the verification semantics independently of any ZK framework, serving as a ground truth reference. Following prior work (Yuan et al., 2023; Siddiq et al., 2023), we then instruct an LLM to generate two Python scripts for test case generation: one producing input-solution pairs $(I,S)$ that should be accepted, and the other producing pairs that should be rejected. This dual design is crucial, as existing code-generation benchmarks provide only accepting examples (verifying acceptance of correct input-solution pairs) but do not evaluate rejection behavior. This ensures that generated programs are tested both for soundness (rejecting invalid ones) and completeness (accepting valid ones).
To ensure the benchmark offers a rigorous and faithful assessment, all test cases are first executed against the Python oracle and then validated and cross-checked by authors for clarity, correctness, and corner-case coverage, following practices in prior benchmarks (Liu et al., 2025; Zhang et al., 2024b; Wang et al., 2022a). Whenever inconsistencies are detected at any stage, the corresponding step is re-executed or manually corrected. If corner-case is found lacking, the test case generator is revised accordingly.
3.2. Measurement Settings
Using ZK-Eval, we measure LLMs’ competencies on ZK code generation. Our study is guided by the following three research questions (RQs):
  • RQ1: Language & Toolchain Knowledge. Do LLMs understand the syntax, semantics, and standard libraries of Circom and Noir?
  • RQ2: Gadget-Level Competence. Can LLMs reliably use or implement constraint gadgets, the building blocks of ZK programs?
  • RQ3: End-to-End Generation. Can LLMs integrate these skills to produce complete, correct ZK programs?
Each RQ corresponds directly to one component of ZK-Eval. For RQ1, models are tested on curated multiple-choice questions covering language knowledge. For RQ2, they are instructed to implement gadget tasks which are then automatically validated against accepting and rejecting test cases. For RQ3, they attempt end-to-end program generation from reformulated HumanEval tasks, with correctness judged by soundness and completeness against dual test suites.
Model Selection. We evaluate 4 representative LLMs: OpenAI’s reasoning models GPT-o4-mini and GPT-o3, along with DeepSeek-V3 and Qwen3. This mix of reasoning and open-source non-reasoning ones allows us to assess both family-specific and general trends. All models are run under identical prompts and budgets, with 10 samples per task to estimate pass rates and reduce variance.
Figure 7. Accuracy of LLMs and Human Experts on the MCQ Benchmark for ZK Language Knowledge.
Figure 8. LLM Performance on End-to-end ZK Benchmarks. Annotated with average accuracy.
Figure 9. Average Error Distribution of Gadget Implementations Across Languages, Types, and Causes.
3.3. (RQ1) Language & Toolchain Knowledge
Fig. 3.2 summarizes performance on the multiple-choice question set. Reasoning models perform best, with GPT-o4-mini reaching 88.1% and GPT-o3 87.2%. Open-source non-reasoning models perform noticeably worse, with DeepSeek V3 at 79.5% and Qwen3 at 78.9%. The human expert for reference scores 88.7%, similar to the reasoning models. Overall, LLMs demonstrate strong competence in ZK language knowledge, achieving accuracy within a few percentage points of the human expert.
Finding ①: On Circom and Noir, LLMs display strong competence in ZK language knowledge. Reasoning models surpass open-source ones and perform similar to human experts.
We further examine performance across question categories and languages. LLMs achieve their highest accuracy on questions about basic syntax and API references, reflecting familiarity with fundamental constructs and widely used library templates. In contrast, accuracy drops on advanced topics such as traits, signal taggings, and generics, which extend expressiveness but are less commonly encountered. The lowest performance is observed on compiler principles, which probe understanding of how source programs are transformed into arithmetic circuits. These results indicate that while LLMs are proficient in surface-level language usage, they struggle with deeper aspects of language design and compilation. Notably, there is no significant difference between Circom and Noir, suggesting that model competence reflects general exposure to ZK languages rather than language-specific familiarity.
Finding ②: LLMs excel at surface-level knowledge such as syntax and API references, but exhibit weaknesses in advanced language features and compiler principles.
These findings are consistent with the nature of the knowledge base: official documents, standard libraries, public tutorials, and open-source examples are all widely accessible training material to LLMs. Therefore, it is anticipated that LLMs can achieve strong performance on such content.
3.4. (RQ2) Gadget-Level Competence
Building on the language knowledge observed in RQ1, we next assess whether LLMs can effectively apply this knowledge in ZK programming, specifically in using or implementing gadgets. Model outputs were analyzed with automated scripts that classified errors based on compiler diagnostics and test outcomes. From their distinct linguistic and functional effects, we grouped errors into six categories: (i) constraint enforcement, where models failed to encode relations using equality constructs ( === in Circom, assert in Noir); (ii) modules, involving incorrect instantiation or usage of library components; (iii) signal handling, specific to Circom, where mistakes occurred in declaring, assigning, or referencing signals; (iv) syntax & parsing, where the generated code violated grammar rules; (v) typing & resolution, where data types or variables were used inconsistently or incorrectly; and (vi) semantic errors, where code compiled but failed functional correctness tests.
Fig. 9 shows the results on the gadget benchmark. Performance is consistently weak: even the best-performing category, logical operations, reached only 52% accuracy in Circom and 49% in Noir. Comparison gadgets scored lower (37% and 20%), while arithmetic and composite gadgets dropped to extremely low accuracy. This contrasts sharply with the strong results on language-level knowledge: when asked to operationalize this knowledge into gadgets, LLMs largely fail to produce correct, testable implementations. This gap reflects the difficulty of code generation for low-resource DSLs (Joel et al., 2024), exacerbated in ZK programming by limited and fragmented resources (Sheybani et al., 2025).
Finding ③: On Circom and Noir languages, LLMs perform consistently weak on using or implementing constraint gadgets across all categories.
Typing and resolution errors are the most frequent. Noir enforces strict type matching in binary operations, disallowing implicit casts between fields and integers, while Circom distinguishes sharply between variables and signals. Models frequently mix types, reference undeclared identifiers, or redefine variables, causing compilation failures. Signal handling errors, unique to Circom, are the next most common, including declaring signals inside control-flow, reassigning them multiple times, or using them before initialization. Constraint enforcement errors also recur: instead of encoding equalities with === or assert, models generate invalid forms such as non-quadratic constraints in Circom or deprecated constructs in Noir. These issues stem from ZK-specific semantics absent in conventional languages.
Performance further varies by gadget type, reflecting library support and algebraic complexity. Logical operations and comparisons, being boolean primitives or supported by libraries, are most tractable. Arithmetic and composite gadgets, such as exponentiation, modulo, floor division, or composites like sum, product, and distinct, lack native support and must be implemented from scratch. These tasks demand reasoning over field arithmetic, careful handling of inverses and negatives, and explicit soundness constraints. The very low accuracy here underscores the inability of current models to generate gadgets beyond those directly provided in libraries.
Finding ④: LLMs handle boolean and comparison gadgets better than arithmetic and composite ones, where ZK-specific semantics and limited library support demand non-trivial reasoning.
3.5. (RQ3) End-to-End Generation
Fig. 8 summarizes the results on the end-to-end generation benchmark. Compared to language knowledge and gadget-level tasks, performance drops substantially across all models. For Circom, accuracy is very low: GPT-o3 reaches only 20% semantically correct and 22% syntactically valid programs, while Qwen3 drops to 5% and 9%. Noir results are somewhat higher but still limited, with GPT-o4-mini at 32% and 40% and Qwen3 at 10% and 19%. Across all settings, semantic correctness lags behind syntactic validity, showing that many compilable programs still fail to encode the intended verification relation.
Several trends emerge from these results. First, Noir tasks consistently yield higher success rates than Circom, likely due to Noir’s simpler type system and cleaner syntax, which reduce opportunities for errors such as signal misuse or type mismatches that are pervasive in Circom. Second, reasoning models substantially outperform open-source ones: GPT-o4-mini and GPT-o3 are 2–3 times more successful than DeepSeek V3 and Qwen3 across both languages, confirming the gap already observed in earlier tasks. Finally, the gap between syntactic and semantic correctness underscores that compilation alone is insufficient for assessing ZK program quality: many outputs are compilable but wrong.
Finding ⑤: LLMs perform poorly on end-to-end code generation: reasoning models beats open-source ones; Noir is easier than Circom; many programs compile but remain incorrect.
4. Augmenting LLMs for ZK Code Generation
Our study reveals a gap between LLMs’ surface-level knowledge of ZK DSLs and their ability to construct correct gadgets or complete programs. Models can recall syntax and library references but struggle to translate this knowledge into valid constraint systems. To bridge this gap, we introduce ZK-Coder, an agentic framework for end-to-end ZK code generation. We first summarize the design principles derived from our findings, then present the system.
4.1. Design Principles Learned from the Findings
Findings ① and ② show that LLMs possess a strong knowledge of ZK DSL syntax, semantics, and standard libraries, achieving near expert-level accuracy even on advanced features. This reflects their absorption of publicly available documentation and indicates their potential for ZK code generation. Despite this, limitations remain that motivate our further exploration.
In Findings ③ and ④, we observe that this knowledge does not translate into reliable gadget-level generation. Generated code frequently suffers from typing errors, signal mismanagement, and faulty constraint enforcement, reflecting the scarcity of training examples with concrete gadget implementations. To bridge this gap between knowing the language and encoding correct programs, we introduce a sketch language in ZK-Coder that abstracts ZKP’s underlying constraint system. To further link them to curated hints, ZK-Coder leverages retrieval-augmented generation (RAG) (Gao et al., 2023) to guide gadget usage. Loosely speaking, RAG trades brittle memorization for grounding: once the needed gadgets are identified, retrieval supplies exact usage patterns that collapse the search space and curb gadget hallucinations. This strategy explicitly mirrors how developers consult a codebook and enables composition of verified snippets for rare gadgets that pretraining scarcely covers.
Finally, Findings ③ and ⑤ highlight end-to-end program generation as the most challenging setting, where generated programs often fail due to syntactic errors or semantic mismatches. These failures directly undermine the reliability and necessitate mechanisms to improve validity and faithfulness. To mitigate these failures, ZK-Coder introduces a correctness-centric interactive loop. The LLM proposes a program, queries a compiler as an oracle to check syntactic validity, and, once compilable, executes test cases as an oracle to check intended semantics. By observing diagnostics and counterexamples, ZK-Coder repairs its code and iterates the “generate-compile-test-repair” loop until both syntactic and semantic checks pass. This interaction transforms failures into actionable feedback and guides candidates toward valid, faithful programs.
4.2. ZK-Coder
Based on these design principles, we developed ZK-Coder, illustrated in Fig. 10. ZK-Coder augments LLMs with a structured pipeline that bridges natural-language problem descriptions and executable ZK programs. It consists of three main stages: ➀ constraint formulation, ➁ constraint-guided analysis and retrieval, and ➂ interactive generation and repair. In ➀, problem statements are translated into lightweight sketch language. Then, in ➁, the sketch is analyzed to extract its constraints, which are used to retrieve relevant usage patterns on gadgets from a curated knowledge base. Finally, in ➂, the LLM generates the target ZK code using the retrieved hints, interactively repairing it with the help of syntax and semantic oracles until correctness is achieved. We detail the design of these stages in this section, and present an illustrative workflow with concrete code examples in the Supplementary Material.
Figure 10. Illustrative Overview on the Design of ZK-Coder.
4.2.1. Constraint Formulation in Sketch Language.
$\begin{array}[]{rrcl}\textsc{Program}&P&::=&\texttt{def}\ \texttt{verify}(\texttt{engine},\ x_{1}\!:\!\tau_{1},\ \ldots,\ x_{n}\!:\!\tau_{n}):\ S^{*}\\[2.0pt] \textsc{Type}&\tau&::=&\texttt{Field}\mid\texttt{Bool}\mid\texttt{Vec[}\tau\texttt{]}\\[2.0pt] \textsc{Stmt}&S&::=&x=E\mid\texttt{engine.add}(\Phi)\mid\texttt{for}\ x\ \texttt{in}\ R:\ S^{*}\mid\texttt{if}\ \Phi:\ S^{*}[\texttt{else}:\ S^{*}]\mid\texttt{return engine}\\[2.0pt] \textsc{Range}&R&::=&\texttt{range}(E)\mid\texttt{range}(E,E)\\[2.0pt] \textsc{Formula}&\Phi&::=&A_{0}\mid\texttt{not}\ \Phi\mid\Phi\ \texttt{and}\ \Phi\mid\Phi\ \texttt{or}\ \Phi\mid\texttt{all}(\Phi\ \texttt{for}\ x\ \texttt{in}\ R)\mid\texttt{any}(\Phi\ \texttt{for}\ x\ \texttt{in}\ R)\\[2.0pt] \textsc{Atom}&A_{0}&::=&E\ \mathrel{\bowtie}\ E\mid G(E_{1},\ldots,E_{n})\\[2.0pt] \textsc{Expr}&E&::=&n\in\mathbb{Z}\mid\texttt{True}\mid\texttt{False}\mid x\mid(E)\mid E\ +\ E\mid E\ -\ E\mid E\ *\ E\mid E\ /\ E\mid E\%E\mid E//E\mid\ldots\\[2.0pt] \textsc{RelOp}&\bowtie&::=&\texttt{==}\mid\texttt{!=}\mid\texttt{<}\mid\texttt{<=}\mid\texttt{>}\mid\texttt{>=}\mid\texttt{and}\mid\texttt{or}\mid\texttt{xor}\\[2.0pt] \textsc{Primitives}&G&::=&\texttt{distinct}(E)\mid\texttt{select}(E,E,E)\mid\texttt{abs}(E)\mid\texttt{sum}(E\ \texttt{for}\ x\ \texttt{in}\ R)\mid\ldots\end{array}$
Figure 11. The Abstract Grammar of the ZK Sketch Language.
The first stage of ZK-Coder reasons and translates a natural-language problem description into a structured representation of ZKP programs. In accordance to our observations in Sec. 3.4 as well as prior work on sketching and planning for code generation (Wang et al., 2024; Deng et al., 2024; Li et al., 2023; Jiang et al., 2024a; Wen et al., 2024a), we introduce a lightweight intermediate language, the ZK Sketch Language (ZKSL; Fig. 11) that shares a similar essence with Python and Z3 ecosystem while abstracting away ZK-specific details. In ZKSL, a Program consists of a single verify function with typed parameters and a sequence of statements (Stmt); statements include assignments, assertions via engine.add($\Phi$), bounded loops and conditionals, and an optional return. Formulas $\Phi$ are built from relational atoms $E\bowtie E$ and gadget calls $G(\cdot)$. Its primitives (literals, arithmetic operators, etc.) align with the gadgets used in Sec. 3.1.1.
ZK-Coder prompts LLMs to generate ZKSL sketches from problem specifications. Then, to ensure well-formedness, ZK-Coder incorporates a checker that validates sketches on syntactic validity, type/arity consistency, and closure. If a sketch is invalid, feedback is provided to the LLM to refine its output. Since ZKSL is ZK-agnostic, it isolates constraint reasoning from ZK-specific syntax and semantics and reduces the generation complexity.
4.2.2. Constraint-Guided Analysis and Retrieval.
In addition to sketching and planning with ZKSL, RAG has been shown powerful to aid LLM-based code generation in low-resource, domain-specific languages (Joel et al., 2024). After obtaining a ZK-agnostic sketch, ZK-Coder analyzes its constraints and retrieves ZK-specific implementation guidance, bridging the abstract ZKSL specification and the concrete gadget implementations required in particular ZK DSLs.
We traverse the ZKSL sketch, collecting each constraint introduced via engine.add($\cdot$). Operators are canonicalized: aliases are normalized (e.g., >= is normalized to GreaterEqThan) and operator variants are distinguished by arity and types (e.g., XOr: Bool $\times$ Bool $\rightarrow$ Bool vs. XOr: Field $\times$ Field $\rightarrow$ Field). Non-constraint constructs (e.g., folded constants, loop statements or conditional guards that emit no constraints) are discarded. The result is an ordered, typed set of constraint primitives corresponding to the required gadgets.
To build a knowledge base for retrieval, we curate a collection of gadget and their usage examples either from the standard libraries of ZK DSLs or from verified implementations. With the knowledge base, ZK-Coder performs constraint-guided retrieval to fetch relevant usage patterns on gadgets for each constraint primitive. Retrieval enforces exact operator/type/arity matching, and hints are emitted in the ZKSL order to preserve program flow during generation. In this way, ZK-Coder effectively grounds generation in precise constraint analysis and structured retrieval.
4.2.3. Interactive Generation and Repair.
With the sketch generated in ➀ and the hints retrieved in ➁, ZK-Coder proceeds to the final stage of end-to-end code generation. As pointed out in prior work (Joel et al., 2024; Wang et al., 2022b; Bi et al., 2024; Skreta et al., 2023), introducing a feedback loop that leverages compilers and test executors as oracles can substantially improve code generation reliability. Motivated by this, ZK-Coder implements an interactive generate-repair loop to ensure both syntactic validity and semantic correctness of the generated ZK programs. The loop alternates between code generation, compilation checks, and execution-based testing, iteratively refining the program until it meets all correctness criteria.
Given the ZKSL sketch specification and constraint-specific hints, the LLM produces a candidate program. Then, ZK-Coder submits the program to the ZKP compiler (i.e., Circom or Noir) for compilation. If ZK-Coder observes compilation errors (e.g., syntax errors, type mismatches, or illegal signal operations), it extracts the diagnostics and feeds them back to the LLM for repair. The LLM refines the program based on this feedback, producing a new candidate that is again submitted for compilation. This process repeats until a compilable program is obtained. With a compilable program, ZK-Coder generates a set of test cases and validates the program on these cases on the fly. If any test case fails, ZK-Coder collects the counterexample and feeds it back to the LLM for further repair. This loop continues until the program successfully passes all test cases or the retry limit is reached. This interactive approach in ZK-Coder converts failures into constructive feedback and augments the LLM’s ability to produce reliable ZK programs.
5. Evaluation
We evaluate ZK-Coder on the task of generating zero-knowledge programs from natural language specifications. Our evaluation is structured around the following research questions, with a focus on end-to-end generation. To clarify, we do not re-evaluate language knowledge (RQ1) or gadget-level coding (RQ2): the former is established in Sec. 3.3, and the latter is handled by our retrieval mechanism in Sec. 4.2.2. Given that said, we include an ablation study RQ5.
  • RQ4 (End-to-End Performance): How does ZK-Coder perform compared to baseline methods in generating correct ZK programs?
  • RQ5 (Ablation Study): What is the effect of each design component, including the ZK sketch language, constraint-guided retrieval, and the interactive repair loop?
  • RQ6 (Robust Generalization): Beyond the standard benchmark, adapted HumanEval, how well does ZK-Coder transfer to newer and more challenging tasks?
  • RQ7 (Failure Analysis): When ZK-Coder fails, what are the common failure patterns, and how might they inform future research?
5.1. Experimental Setup
Our experiments are organized according to the research questions outlined above.
For RQ4 (End-to-End Performance), we evaluate on the adapted HumanEval benchmark from ZK-Eval, using the two strongest OpenAI’s reasoning models from our measurement study: GPT-o4-mini and GPT-o3. This comparison establishes how well ZK-Coder outperforms baseline methods in generating correct and provable ZK programs.
For RQ5 (Ablation Study), we conduct an ablation study on GPT-o4-mini, selectively disabling each design component of ZK-Coder, including the sketch language, constraint-guided retrieval, and the interactive repair loop. This aims to quantify their individual impact on overall performance.
For RQ6 (Robust Generalization), we construct a new dataset following the adaptation methodology of ZK-Eval using LiveCodeBench. Unlike the simple and potentially contaminated HumanEval, this benchmark contains harder, contamination-free, and more recent tasks. We evaluate whether ZK-Coder continues to maintain high performance under more challenging scenarios.
Finally, for RQ7 (Failure Analysis), we analyze cases where ZK-Coder fails, categorizing the dominant error patterns and their underlying causes. This analysis highlights limitations of the current pipeline and suggests avenues for future research.
All experiments are conducted under a fixed repair budget of up to 8 iterations for syntax checking and repair, and 1 iteration for semantic checking and comparison. Each task is executed 10 times to compute Pass@1 accuracy, ensuring statistical stability. The evaluation is conducted on a Linux (Ubuntu 22.04 LTS) server equipped with 256 GiB of RAM.
5.2. RQ4: End-to-End Performance
We evaluate the end-to-end performance of ZK-Coder. To the best of our knowledge, since no prior work directly addresses zero-knowledge program generation from natural language, we establish a simple but meaningful baseline: directly prompting the model to produce circuit code from the task description and input specification, without any sketching, retrieval, or repair. This setting captures the raw capability of LLMs and provides a lower bound against which the contributions of ZK-Coder can be measured.
During the evaluation, we meticulously document the performance for the pipeline in three key stages: (1) Sketch correctness measures whether the initial ZK sketch produced from the natural language description is correct by passing all hidden test cases; (2) Repair pass rate is conditioned on all correctly generated sketches, and reports the fraction of these that can be successfully repaired by the interactive repair loop; (3) Program correctness is conditioned on all successfully repaired programs, and measures the proportion that pass all hidden test cases. The final overall accuracy aggregates across the full pipeline, reflecting end-to-end program generation accuracy. In addition, we report the average token cost for each model, which quantifies the spending.
Table 3. Comparative Per-stage Performance (Accuracy, Tokens) of ZK-Coder and Baseline.
Language Model ZK-Coder Per-stage Performance ZK-Coder Overall Baseline
(1) Sketch (2) Repair (3) Program Accuracy Avg. Accuracy Avg.
Correctness Pass Rate Correctness (Pass@1) Token (Pass@1) Token
Circom GPT-o4-mini 94.12% 91.72% 96.59% 83.38% 2350.31 17.35% 436.98
GPT-o3 97.30% 94.73% 96.82% 89.23% 2253.29 20.29% 433.35
Noir GPT-o4-mini 94.12% 97.97% 97.66% 90.05% 648.95 32.21% 152.18
GPT-o3 97.21% 97.43% 99.38% 94.12% 765.46 27.94% 195.34
Table 3 summarizes the comparative performance of ZK-Coder and the baseline across both Circom and Noir. Across all settings, ZK-Coder achieves substantial gains over the baseline. On Circom, GPT-o4-mini and GPT-o3 reach overall pass rates of 83.38% and 89.23%, respectively, compared to only 17.35% and 20.29% for the baseline. Noir shows a similar trend, with GPT-o4-mini reaching 90.05% and GPT-o3 achieving the highest overall rate of 94.12%, far outperforming the corresponding baseline results of 32.21% and 27.94%.
The per-stage breakdown highlights the robustness of the pipeline. Sketch correctness exceeds 94% across all configurations, showing that initial translation into ZK sketches is highly reliable. The repair loop maintains strong success rates, above 91% in Circom and above 97% in Noir, ensuring that most sketches can be generated into valid programs. Final program correctness surpasses 96% in all cases, confirming that the integration of sketching, repair, and semantic validation yields consistently executable and provable programs.
We also note that ZK-Coder incurs a higher token cost compared to the baseline, which is expected since our agentic pipeline includes sketching, RAG and interactive repair. However, the cost remains modest in practice, with each task requiring fewer than 3,000 tokens on average, corresponding to less than 0.1 USD per task, making the gains in correctness highly cost-effective.
Answer to RQ4: ZK-Coder delivers a significant improvement in end-to-end performance over the baseline method. These results confirm that our pipeline substantially enhances the performance on LLM-generating ZK programs.
5.3. RQ5: Ablation Study
Figure 12. Accuracy Comparison of ZK-Coder Ablation Variants.
To assess the role of each major component in ZK-Coder, we conduct an ablation study where individual modules are disabled or replaced. The following variants are considered:
  • No RAG (V1): Disables constraint-guided retrieval augmented generation, forcing the model to rely solely on its internal knowledge.
  • No Sketch (V2): Removes the intermediate ZK sketch reasoning and directly prompts the model to output full programs with RAG and the interactive repair loop.
  • No Syntax Repair (V3): Eliminates the interactive repair loop for syntax errors.
  • No Semantic Repair (V4): Retains syntax repair but disables semantic repair.
  • Similarity-based RAG (V5): Replaces our constraint-guided retrieval with a naive similarity-based retrieval using surface-level embedding similarity.
  • Only Repair (V6): Omits both retrieval and sketching, relying exclusively on the repair loop applied to direct generations.
Figure 12 presents the pass@1 accuracy for each ablation variant across Circom and Noir. The results show that every component plays a critical role in the overall success of ZK-Coder. Removing RAG (V1) reduces accuracy from 83.38% to 67.21% on Circom, and from 90.05% to 89.22% on Noir. Disabling the sketch reasoning stage (V2) causes further degradation (71.32% on Circom, 87.50% on Noir), highlighting the value of sketching as a scaffold for task understanding and constraint reasoning. The most severe drop arises from removing syntax repair (V3), with accuracy falling to 35.00% on Circom and 41.39% on Noir, demonstrating that interactive repair is indispensable. Disabling semantic repair (V4) also leads to measurable reductions (81.18% on Circom, 87.99% on Noir), confirming that semantic validation contributes to robustness. Substituting constraint-guided retrieval with a naive similarity-based strategy (V5) yields weaker performance (69.87% on Circom, 89.41% on Noir), suggesting that constraint-guided retrieval is more effective than surface-level matching. Finally, the “Only Repair” variant (V6) performs poorly, especially on Circom (59.12%), indicating that repair alone cannot compensate for the absence of sketching and retrieval.
Answer to RQ5: The ablation study shows that the strong performance of ZK-Coder relies on the synergy of sketching, constraint-guided RAG, and interactive repair. Removing or weakening any of these components causes substantial accuracy drops.
5.4. RQ6: Robust Generalization
To assess whether ZK-Coder sustains strong performance on harder, more recent, and contamination-free tasks, we evaluate on LiveCodeBench (Jain et al., 2024), a recently introduced benchmark designed to be contamination free. We select all tasks labeled as medium difficulty, filter and adapt them using the methodology introduced in Sec. 3.1.3, resulting in a set of 34 tasks (details refer to the Supplementary Material). This offers a more challenging and realistic testbed for ZK program generation.
Table 4. ZK-Coder Code Generation Performance on the Adapted LiveCodeBench Benchmark.
Language ZK-Coder Baseline
Circom 72.25% 2.97%
Language ZK-Coder Baseline
Noir 82.32% 14.72%
Table 4 presents the results. ZK-Coder achieves 72.25% accuracy on Circom and 82.32% on Noir, dramatically outperforming the baseline scores of 2.97% and 14.72%, respectively. These results indicate that even on a harder and contamination-free benchmark, ZK-Coder continues to deliver strong accuracy and robust generalization.
Answer to RQ6: ZK-Coder’s performance on LiveCodeBench confirms that the pipeline generalizes effectively to harder and contamination-free tasks.
5.5. RQ7: Failure Analysis
Table 5. Distribution of ZK-Coder’s Code Generation Failures.
Language Benchmark Repair Budget Sketch False False Mixed False
Exceed Incorrect Accept Reject Accept & Reject
Circom Humaneval 46.09% 34.78% 0.87% 14.78% 3.48%
LiveCodeBench 36.26% 37.36% 9.89% 15.38% 1.10%
Noir Humaneval 17.33% 53.33% 5.33% 17.33% 6.67%
LiveCodeBench 23.53% 58.82% 0.00% 17.65% 0.00%
Table 5 reports the distribution of ZK-Coder’s failure cases across Circom and Noir on both HumanEval and LiveCodeBench. Each column corresponds to a distinct failure mode. Repair Budget Exceed indicates cases where the repair budget was exhausted without producing any compilable program. Sketch Incorrect denotes that the generated ZKSL sketch failed hidden test cases, suggesting a misunderstanding of the task requirements. False Accept corresponds to programs that erroneously accept invalid witnesses, while False Reject refers to programs that incorrectly reject valid witnesses. Mixed False Accept & Reject captures programs that simultaneously exhibit both types of errors, reflecting severe constraint mis-specification.
The data reveal that the two dominant sources of failure are Repair Budget Exceed and Sketch Incorrect. In both Circom and Noir, these two together account for over 80% of failures on both benchmarks. Repair Budget Exceed often arises when the model is unable to generate any compilable program, which due to gaps in pretraining exposure to ZK-specific constructs. Sketch Incorrect, on the other hand, point to deeper misunderstandings of the natural language description, where the task is not properly captured in the ZKSL representation. The less frequent categories (false accept, false reject, and mixed errors) typically stem from incorrect constraint enforcement when translating a valid sketch into a concrete circuit. These errors indicate local implementation mistakes rather than systemic failures of the pipeline.
Answer to RQ7: In summary, Repair Budget Exceed and Sketch Incorrect constitute the most critical bottlenecks for robust end-to-end generation. The former highlights the need for more ZK-specific training data to improve raw generation capabilities, while the latter underscores the difficulty of grounding natural language descriptions in precise constraints.
6. Related Work
LLM for DSL Code Generation.
Emerging research has investigated the use of LLMs for generating code in DSLs, a task made challenging by limited training data and specialized syntax (Joel et al., 2024). To mitigate these difficulties, grammar prompting has been proposed to guide LLMs with partial grammars and thereby improve the well-formedness of generated programs (Wang et al., 2023). RAG techniques further enhance DSL generation by supplying relevant API documentation or code samples (Pimparkhede et al., 2024; Bassamzadeh and Methani, 2024). For DSLs where correctness guarantees are critical, approaches based on formal verification (Councilman et al., 2025) and verified lifting (Bhatia et al., 2024) strengthen reliability by coupling generated code with machine-checkable proofs. Our work advances this line of research by focusing on DSLs for zero-knowledge proofs.
Code Generation Benchmarks.
Most benchmarks for LLM-based code generation focus on general-purpose languages (Jain et al., 2024; Chen et al., 2021), with limited support for DSLs. A recent survey (Joel et al., 2024) reviews existing DSL benchmarks and emphasizes the lack of standardized evaluation resources, highlighting the need for benchmarks that capture DSL-specific paradigms, idioms, and practical applications. In the case of ZKPs, to our knowledge, no systematic benchmark exists for ZKP code generation that reflects ZK-specific requirements, namely proving a computation, a static arithmetic-circuit model, and correctness defined by both acceptance and rejection. We provide a solution for this gap with ZK-Eval, studying code generation bottlenecks and evaluating end-to-end tasks.
Software Engineering for ZKP
The software engineering community has recently begun to explore tools and methodologies for ZKP development. Numerous researches have proposed to improve the security of ZKPs. Static analysis tools such as Circomspect (Dahlgren, 2022), ZKAP (Wen et al., 2024b), QED2 (Pailoor et al., 2023), and halo2-analyzer (Soureshjani et al., 2023) identify specific vulnerability patterns, while testing tools like MTZK (Xiao et al., 2025) and zkFuzz (Takahashi et al., 2025) use fuzzing to expose security issues. In this work, we focus on reducing the burden of writing ZKP programs by leveraging LLMs’ capabilities, complementing these efforts by helping developers audit and refine their code.
7. Discussion, Limitation and Future Work
ZKP Program Complexity.
Program complexity for ZKPs is measured by the size of arithmetic circuits. It is a critical factor in the efficiency and scalability of ZKP systems, directly determining prover cost and overall performance. To reduce that, a range of compilers have been proposed to optimize circuit construction (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022), while new protocols reduce complexity through advanced arithmetizations with custom gates and lookup arguments (Chen et al., 2023; Gabizon et al., 2019). On the programmer side, several techniques further improve efficiency, such as adopting ZK-friendly hash functions (Grassi et al., 2021), gate packing, and aggregating bits into single field elements. For LLM-based code generation, however, producing efficient circuits is more difficult than generating correct ones. As future work, we plan to survey optimization techniques and guide LLMs in incorporating these strategies.
LLM Training and Fine-tuning.
Recent studies show that LLMs struggle with DSLs and other low-resource languages due to the lack of high-quality training data (Joel et al., 2024), a challenge that is particularly acute in the context of ZKPs (Sheybani et al., 2025). This scarcity also limits the feasibility of fine-tuning approaches. Our work takes a step toward mitigating this limitation by systematizing the generation of program examples in ZK DSLs, which can support targeted training and domain adaptation of LLMs. These examples may be used as a complementary resource for future research in ZK code generation.
Threat to Validity.
Our study on language knowledge relies on multiple-choice QA pairs that were initially generated with LLM assistance. Although some of these questions may contain leading hints or incorrect answers, we address this risk through repeated human validation and cross checking, ensuring that any residual noise has limited influence on the results. The evaluation in ZK-Eval employs carefully designed test cases that check both accepting and rejecting behaviors of generated programs and gadget usages. While test cases cannot fully guarantee correctness, they are constructed to include representative corner cases and to enforce soundness and completeness requirements, thereby reducing the likelihood of false positives. Furthermore, the gadget suite used by ZK-Eval and ZK-Coder consists of 35 representative components selected for their prevalence in ZK libraries and for their coverage of core patterns. This set strikes a balance between complexity and tractability, spanning arithmetic, boolean, comparison, etc., and has already demonstrated clear performance improvements in evaluation. Looking forward, the suite is designed for incremental extension, allowing more advanced and cryptographic composites to be incorporated in future work.
8. Conclusion
We presented ZK-Eval, the first benchmark for systematically assessing LLM capability in ZK program synthesis, and showed that while models master language knowledge, they struggle with gadget construction and end-to-end synthesis. To address these gaps, we introduced ZK-Coder, a retrieval-augmented framework that guides LLMs from abstract constraint sketches to executable ZK code through constraint-guided retrieval and syntax-semantic repair. We show that ZK-Coder significantly improves the code generation capability of LLMs on both Circom and Noir.
Appendix A Illustrative Workflow Example
We demonstrate the workflow using ZK-Coder to generate a ZKP program that can be used to prove the correctness of a Sudoku assignment in Fig. 13. Given a $4\times 4$ Sudoku grid, the goal is to verify that each row, column, and $2\times 2$ block contains all digits 1144 exactly once.
Figure 13. ZK-Coder’s Illustrative Example Workflow on Proving Sudoku Correctness.
In this illustrative workflow example, ZK-Coder first reads the problem description and formulates it in ZKSL, encoding the necessary constraints on the Sudoku grid. The checker validates the sketch and analyzes it to extract the set of constraint gadgets required. For each gadget, relevant implementation hints are retrieved, and an initial Circom implementation is generated. The compiler then checks this program for syntactic validity, returning diagnostics that guide interactive refinement. Once the code compiles, the model is prompted to generate a small test instance. If the resulting prover behavior deviates from expectations, the model is asked to repair the program. This repair loop continues until either the program passes the LLM-supplied tests or the retry budget is exhausted.
We also present two concrete program examples generated by ZK-Coder for this Sudoku problem, one in Circom (Fig. 14) and one in Noir (Fig. 15). Both programs implement the same verification logic. We annotated the source code with comments to point out key sections, such as input signal declarations, gadget imports or implementations, and composing gadgets with field arithmetic operations to implement the constraints outlined in the ZKSL sketch.
Figure 14. Example Circom Program Generated by ZK-Coder.
Figure 15. Example Noir Program Generated by ZK-Coder.
Appendix B Filtered Tasks in Dataset Curation
As introduced in the main paper, our dataset construction is based on adapting existing programming tasks from HumanEval (Chen et al., 2021) and LiveCodeBench (Jain et al., 2024). To ensure that tasks are suitable for ZK DSLs (no floating-point numbers, no string manipulation, no I/O, deterministic output length, etc.), we filter the tasks using a combination of LLM assisted filtering and manual inspection. We now present the complete set of filtered tasks in Table 6 and Table 7.
Table 6. Filtered tasks from HumanEval.
Filtered Task IDs from HumanEval (Total 68 Tasks)
3 5 8 9 13 24 31 33 35 36 37 39 40 41 42 43 46
49 52 53 55 57 59 60 62 63 65 69 70 72 73 75 76 77
83 85 88 90 94 97 102 106 108 109 110 114 115 116 121 122 123
126 127 128 131 135 138 139 142 145 146 147 150 151 152 155 157 159
Table 7. Filtered tasks from LiveCodeBench.
Filtered Task IDs from LiveCodeBench (Total 34 Tasks)
43 44 62 66 78 94 119 124 129 153 156 172 203 225 232 240 243
250 261 265 280 290 297 305 322 323 336 346 353 367 373 375 384 389
Appendix C Supplementary Results
In this section, we present supplementary results that complement the main findings of the paper but were omitted earlier due to space limitations.
C.1. Error Distribution For Gadget Implementation
Previously, we reported the error distribution averaged across four models. Here, we provide the per-model error distributions. The breakdown by individual model is consistent with the overall trend: OpenAI’s reasoning models outperform open-source models, and the most common error types remain consistent across models.
Figure 16. Error Distribution on GPT-o4-mini.
Figure 17. Error Distribution on GPT-o3.
Figure 18. Error Distribution on DeepSeek-V3.
Figure 19. Error Distribution on Qwen3.
Appendix D Complete Set of Constraint Gadgets
The constraint gadgets used in the main paper during both measurement and augmentation. In measurement, they serve as the targets for evaluating LLMs’ ability to correctly employ or implement gadgets when expressing constraints. In augmentation, they are treated as primitive operators in the sketch language, and in constraint-guided retrieval we prepared knowledge snippets for each gadget to assist LLMs in code generation.
The collection of gadgets is consolidated from three sources: (i) the official Circom standard library (iden3, 2025), (ii) the ZK-kit library (Nico Serrano, [n. d.]), and (iii) the Z3Py API (Research, [n. d.]). The complete list is shown in Table 8. We classify gadgets into five categories: Logical, Bitwise, Comparison, Arithmetic, and Composite, according to their functionality. The typing column specifies input and output types, which are used to disambiguate operators during constraint-guided retrieval. While this set is not exhaustive, it spans a wide range of common constraints in ZK programming. We expect to extend it with additional gadgets in future work.
Table 8. List of constraint gadgets used in measurement and augmentation.
Name Category Typing Source
And Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
And Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Or Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Or Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib,Z3Py
XOr Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
XOr Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Not Logical Bool $\rightarrow$ Bool Circomlib, Z3Py
Or Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
Or⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
And Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
And⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
XOr Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
XOr⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
Equal Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
NotEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
LessThan Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
LessThanOrEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
GreaterThan Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
GreaterThanOrEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
Add Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Subtract Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Multiply Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Divide Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
Modulo Arithmetic Field $\times$ Field $\rightarrow$ Field ZK-Kit, Z3Py
Power Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
FloorDivide Arithmetic Field $\times$ Field $\rightarrow$ Field ZK-Kit, Z3Py
Negate Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Absolute Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Inverse Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Sign Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
Conditional Composite Bool $\times$ Field $\times$ Field $\rightarrow$ Field Circomlib, Z3Py
Conditional Composite Bool $\times$ Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
Sum Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
Product Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
Distinct Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py

### v2

@@
 Abstract.
-Zero-knowledge proofs (ZKPs) are increasingly deployed in domains such as privacy-preserving authentication, blockchain scalability, and secure finance.
-However, authoring ZK programs remains challenging: unlike mainstream programming, ZK development requires reasoning about finite field arithmetic, constraint systems, and gadgets, making it knowledge-intensive and error-prone.
-While large language models (LLMs) have demonstrated strong code generation capabilities in general-purpose languages, their effectiveness for ZK programming, where correctness hinges on both language mastery and gadget-level reasoning, remains unexplored.
-To address this gap, we propose ZK-Eval, a domain-specific evaluation pipeline that probes LLM capabilities at three levels: language knowledge, gadget competence, and end-to-end program generation.
-Our evaluation of four state-of-the-art LLMs reveals that models excel at surface-level syntax but struggle with gadget usage and semantic correctness, often yielding incorrect programs.
+Zero-knowledge proofs (ZKPs) are increasingly deployed in domains such as privacy-preserving authentication, verifiable computation, and secure finance.
+However, authoring ZK programs remains challenging: unlike conventional software development, ZK programming manifests a fundamental paradigm shift from imperative computation to declarative verification.
+This process requires rigorous reasoning about finite field arithmetic and complex constraint systems (which is rare in common imperative languages), making it knowledge-intensive and error-prone.
+While large language models (LLMs) have demonstrated strong code generation capabilities in general-purpose languages, their effectiveness for ZK programming, where correctness hinges on both language mastery and constraint-level reasoning, remains unexplored.
+To address this gap, we propose ZK-Eval, a domain-specific evaluation pipeline that probes LLM capabilities on ZK programming at three levels: language knowledge, algebraic primitive competence, and end-to-end program generation.
+Our evaluation of four state-of-the-art LLMs reveals that while models demonstrate strong proficiency in language syntax, they struggle when implementing and composing algebraic primitives to specify correct constraint systems, frequently producing incorrect programs.
 Based on these insights, we introduce ZK-Coder, an agentic framework that augments LLMs with constraint sketching, guided retrieval, and interactive repair.
-Experiments on Circom and Noir show substantial gains, with success rates improving from 17.35% to 83.38% and from 32.21% to 90.05%, respectively.
-With ZK-Eval and ZK-Coder, we establish a foundation for systematically measuring and augmenting LLMs in ZK code generation to lower barriers for practitioners and advance trustworthy computation.
-Zero-Knowledge Proofs, Large Language Models, Code Generation
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournalvolume: Xjournalnumber: Xarticle: Xjournalyear: 2025publicationmonth: 9copyright: noneccs: Software and its engineering Automatic programmingccs: Security and privacy Cryptographyccs: Software and its engineering Domain specific languagesccs: Computing methodologies Natural language processing
+Experiments with GPT-o3 on Circom and Noir show substantial gains, with success rates improving from 20.29% to 87.85% and from 28.38% to 97.79%, respectively.
+With ZK-Eval and ZK-Coder, we establish a new basis for systematically measuring and augmenting LLMs in ZK code generation to lower barriers for practitioners and advance privacy computing.
+copyright: noneccs: Software and its engineering Automatic programmingccs: Security and privacy Cryptographyccs: Software and its engineering Domain specific languagesccs: Computing methodologies Natural language processing
 1. Introduction
@@
 In practice, this means proving that a computation or claim is correct while keeping sensitive inputs private.
-ZKPs have moved far beyond theory: they are now central to privacy-preserving authentication, blockchain scalability, and emerging applications across finance and security domains (Gupta, 2025; Lavin et al., 2024; team, 2024).
+ZKPs have moved far beyond theory: they are now central to privacy-preserving authentication, blockchain scalability, verifiable computation, and emerging applications across finance and security domains (Sun et al., 2024; Gupta, 2025; Lavin et al., 2024; team, 2024).
 As demand for trustworthy computation grows, the ability to author, test, and verify ZKP programs has become a first-class concern in the software engineering community (Takahashi et al., 2025; Hochrainer et al., 2024; Xu et al., 2025; Pailoor et al., 2023; Jiang et al., 2025a).
@@
 Unlike mainstream programming, which builds on the von Neumann model and imperative execution, ZKP programs boil down to specifying and proving mathematical constraints over finite fields.
-Its development requires reasoning about finite field arithmetic, constraint systems, and gadgets — a novel concept advocated by modern ZKP frameworks to encode reusable, domain-specific building blocks.
+In essence, this represents a paradigm shift from computation to verification.
+Its development requires reasoning about finite field arithmetic and constraint systems.
 As shown in Fig.
 1 and Table 1, ZK programs do not describe how to execute a computation but how to prove its correctness by compiling constraints into arithmetic circuits.
-Developers must specify equations over inputs, outputs, and intermediate variables, carefully wiring gadgets to enforce correctness.
-Even small omissions or miswirings can yield satisfiable but unsound constraint systems, leading to errors that compilers rarely flag.
-This makes ZK development knowledge-intensive, error-prone, and inaccessible to non-experts, despite significant progress in proof systems and verification tools (Chaliasos et al., 2024; Pailoor et al., 2023; Takahashi et al., 2025).
-Table 1. Comparison of Mainstream and ZK Programming.
+Developers must specify equations over inputs, outputs, and intermediate variables, carefully wiring algebraic primitives to specify constraints enforcing correctness.
+Even small omissions or miswirings can yield unsound or incomplete constraint systems, leading to errors that compilers rarely flag.
+This makes ZK development knowledge-intensive, error-prone, and inaccessible to non-experts, despite significant progress in proof systems and verification tools (Chaliasos et al., 2024; Pailoor et al., 2023; Takahashi et al., 2025; Kolozyan et al., 2025; Chaliasos et al., 2025).
+Table 1. Mainstream & ZK Program Comparison.
 Aspect Mainstream ZK Program
@@
 Comp. Model von Neumann Arch Arithmetic Circuits
-Purpose to Compute to Prove
+Purpose to Compute to Prove/Verify
 Data Types Int/Float/Str/… Finite Field Elements
 Mutability Dynamic Static
-Turing-complete? Yes No
 Control Flow If/Loop/Recursion Limited
@@
 Answering this question requires more than surface-level benchmarking.
-Compared to programming in mainstream languages, success in ZK development hinges on two competencies: (i) knowledge of domain-specific languages (DSLs) and their toolchains, and (ii) mastery of gadgets as building blocks for constraint systems.
-Existing benchmarks focus only on mainstream languages and overlook these unique requirements.
-To address this gap, we introduce ZK-Eval, a domain-specific evaluation pipeline that probes LLM capability at three levels: language knowledge, gadget-level competence, and end-to-end program code generation (left panel of Fig.
+Compared to programming in mainstream languages, success in ZK development hinges on three competencies: (i) knowledge of domain-specific languages (DSLs) and their toolchains, (ii) mastery of algebraic primitives as atomic building blocks for constraint systems, and (iii) proficiency in modeling and specifying complex verification constraints.
+Existing benchmarks focus only on mainstream programming paradigms and overlook these requirements.
+To measure this knowledge gap, this paper first introduces ZK-Eval, a domain-specific evaluation pipeline that probes LLM capability at three levels: language knowledge, primitive-level competence, and end-to-end program code generation (left panel of Fig.
 2).
-We anticipate that this design would provide a structured way to pinpoint where models succeed, where they fail, and why, rather than treating code generation as a monolithic task.
-Applying ZK-Eval to four state-of-the-art LLMs (including advanced reasoning models like GPT-o4-mini and GPT-o3) reveals a clear pattern.
-While models perform strongly on surface-level language knowledge, their accuracy drops significantly when reasoning about gadgets or assembling complete programs, often producing code that compiles but fails semantically.
-These results underscore both the promise of LLMs in lowering barriers to ZK development and the need for augmentations to bridge the gap between syntactic knowledge and reliable code generation.
+We anticipate that this design would provide a structured way to pinpoint where models succeed, where they fail, and why.
+As will be detailed in § 3, applying ZK-Eval to four SoTA LLMs (including advanced reasoning models like GPT-o4-mini and GPT-o3) reveals a clear pattern.
+While models perform strongly on language knowledge, their accuracy drops significantly (below 30%) when reasoning about primitives or assembling verification constraints, highlighting a critical gap between syntactic knowledge and end-to-end generation.
+These results underscore both the potential of LLMs to democratize ZK development and the need for targeted augmentations to bridge this capability gap.
 Motivated by these findings, we present ZK-Coder (right panel of Fig.
-2), an agentic framework that augments LLMs with constraint sketching, retrieval, and interactive repair.
-By grounding generation in constraint reasoning and guided gadget usage, ZK-Coder achieves substantial gains in reliability, improving success rates from 17.35% to 83.38% on Circom and from 32.21% to 90.05% on Noir.
+2), which combines three key components: (i) a sketch layer (ZKSL) that explicitly models constraints using algebraic primitives, (ii) retrieval-augmented generation to ground specifications in verified implementations, and (iii) interactive refinement to fix constraint misspecifications.
+We observe that ZK-Coder achieves substantial gains, improving code generation success rates from 20.29% to 87.85% on Circom and from 28.38% to 97.79% on Noir (using GPT-o3).
+Ablation studies further confirm that sketching, guided retrieval and interactive refinement are indispensable; removing either component causes a significant performance drop (over 10%).
 In summary, our contributions are as follows:
@@
   3. (3)
-    Zero-Knowledge: If a statement is true, the verifier learns nothing beyond the validity of the statement itself.
-These properties make ZKPs particularly useful in privacy-preserving applications (Gupta, 2025; team, 2024; Lavin et al., 2024).
-However, achieving these properties relies on the correct implementation of the underlying cryptographic primitives and protocols with respect to the high-level intent.
-Consequently, writing ZKP programs can be challenging due to the complexity of the underlying field-arithmetic mathematics and the need for specialized knowledge in cryptography (Pailoor et al., 2023; Wen et al., 2024b; Chaliasos et al., 2024).
-Table 2. Tools for Programming Zero-Knowledge Proofs
-Frontend DSL Compiler GitHub repository
-Circom (Bellés-Muñoz et al., 2022) Circom Rust iden3/circom (1.5K)
-Noir (Aztec Labs, 2025) Rust-like Rust noir-lang/noir (1.2K)
-Leo (Chin et al., 2021) Rust-like Rust AleoHQ/leo (4.8K)
-ZoKrates (Contributors, 2025c) Python-like Rust Zokrates/ZoKrates (1.9K)
-Loosely speaking, developers need to translate their high-level intent into an arithmetic representation $R(x,w)=1$, where $x$ denotes public inputs, $w$ denotes the private witness, and $R$ is a collection of polynomial equations over field elements.
-To ease the development process, several programming frameworks have emerged, providing higher-level abstractions and DSLs for expressing ZKP computations (Sheybani et al., 2025; Ozdemir et al., 2022; Thomas, 2022).
-In Table 2, we review several representative frontend frameworks, including Circom, Noir, and others.
-These frameworks differ not only in syntax but also in how they expose the underlying constraint system, which directly impacts the difficulty of automatic code generation.
-Circom requires developers to think in terms of low-level circuit wiring and explicit constraints, which poses challenges for LLM-based generation since the model must produce correct and efficient circuit templates (Hochrainer et al., 2024; Chaliasos et al., 2025; Takahashi et al., 2025).
-Noir, by contrast, adopts a Rust-like syntax and enforces type safety, making it closer to conventional programming languages and thus more approachable for LLMs trained on general-purpose code corpora (Joel et al., 2024; Jiang et al., 2024b, 2025b).
-As a concurrent effort, there are also emerging frameworks that target general-purpose languages, such as SP1 (Labs, 2025), Nexus (Contributors, 2025a), and RISC Zero (Contributors, 2025b), which compiles Rust code into a restricted instruction set and generates ZK proofs over the execution trace within a virtual machine.
-However, these frameworks often come with significant performance overheads due to the complexity of instruction emulation within the virtual machine.
-Therefore, we consider them as an orthogonal direction to our work for reducing the burden of writing ZKP programs.
-Taken together, these tools illustrate the diverse design space of ZKP programming frontends, and highlight why aligning LLM-based code generation with specific framework characteristics is essential for usability and correctness.
-In this work, we focus on Circom and Noir for two main reasons.
-First, they are the most widely adopted general-purpose ZK frontends in both academia and industry, with active communities and production deployments (e.g., Circom in Ethereum-based rollups, Noir in Aztec).
-This ensures that advances in LLM-assisted code generation for these languages can have immediate practical impact.
-Second, they represent two complementary design paradigms that are highly relevant for evaluating code generation: Circom exposes low-level, circuit-oriented constructs where the challenge for LLMs is to correctly translate high-level intent into constraint wiring, whereas Noir provides a Rust-like syntax with strong typing and higher-level abstractions, which better aligns with the training data of current LLMs.
-By studying both, we capture the spectrum from circuit-level specification to program-level ergonomics, allowing us to analyze how LLMs handle different abstraction levels in ZKP development.
+    Zero-Knowledge: The verifier learns nothing beyond the validity of the statement itself.
+These properties make ZKPs particularly useful in verifiable computation and privacy-preserving applications (Gupta, 2025; team, 2024; Lavin et al., 2024).
+However, ZK programming necessitates a fundamental paradigm shift from imperative computation to declarative verification.
+Rather than writing instructions to calculate a result, developers must model their high-level intent as an arithmetic representation $R(x,w)=1$ over a finite field, where $x$ denotes public inputs and $w$ denotes the private witness.
+This requirement to faithfully enforce logical intent through field-arithmetic constraints is notoriously arduous for humans; even minor errors in modeling can lead to under-constrained circuits (compromising soundness) or over-constrained systems (breaking completeness) (Pailoor et al., 2023; Wen et al., 2024a; Chaliasos et al., 2024; Kolozyan et al., 2025).
+To mitigate this complexity, several programming frameworks have emerged, offering a diverse range of abstractions. We categorize these into two primary design paradigms:
+  1. (1)
+    Constraint-Oriented DSLs.
+Frameworks like Circom and Noir provide specialized syntax for expressing verification relations.
+Circom requires developers to think in terms of low-level circuit wiring and explicit signal management, which poses significant challenges for LLMs tasked with producing correct circuit templates (Hochrainer et al., 2024; Chaliasos et al., 2025).
+Conversely, Noir adopts a Rust-like syntax with strong typing, aligning more closely with the general-purpose coding corpora on which current LLMs are trained (Joel et al., 2024; Jiang et al., 2024a).
+  2. (2)
+    General-Purpose zkVMs.
+Emerging solutions like SP1 (Labs, 2025) and RISC0 (Contributors, 2025) allow developers to compile standard Rust code into a restricted instruction set, generating proofs over the execution trace within a virtual machine.
+While they abstract away the constraint system entirely, they often incur substantial performance overhead.
+We consider these an orthogonal approach to our work, as we focus on the direct generation of circuits.
+In this work, we focus on Circom and Noir because they represent the most widely adopted frontends in both academia (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022; Pailoor et al., 2023; Wen et al., 2024a) and industry (e.g., in Ethereum rollups and Aztec).
+Furthermore, they occupy two complementary points on the abstraction spectrum: Circom tests an LLM’s ability to perform circuit-level specification and constraint wiring, while Noir evaluates how effectively models can adapt conventional coding proficiency to the verification-oriented paradigm.
+By studying both, we capture the core technical challenges of ZKP development while ensuring our findings have immediate practical relevance.
 2.2. LLM-based Code Generation
-Large Language Models (LLMs) have emerged as powerful assistants in software development, enabling code generation from natural language prompts with tools like GitHub Copilot, Cursor, and model-driven agents quickly translating intent into code (Jiang et al., 2024b, 2025b).
+Large Language Models (LLMs) have emerged as powerful assistants in software development, enabling code generation from natural language prompts with tools like GitHub Copilot, Cursor, and model-driven agents quickly translating intent into code (Jiang et al., 2024a, 2025b).
 These systems are reshaping engineering workflows, aiding in writing, testing, and documenting code by drawing on vast amounts of publicly available repositories and documentation (GitHub, 2025).
 Framework.
-Modern LLM-based code generation typically follows an agentic framework (Zhang et al., 2024a; Wang et al., 2025), where the model acts as an intelligent assistant, interpreting user intent, retrieving relevant information, and producing code snippets or entire functions as needed.
-This process often involves multiple iterations, with the model refining its output based on user feedback and additional context.
+Modern LLM-based code generation typically follows an agentic framework (Zhang et al., 2024a; Wang et al., 2025a), where the model acts as an intelligent assistant, interpreting user intent, retrieving relevant information, and producing code snippets or entire functions as needed.
+This process often involves multiple iterations, with the model refining its output based on user feedback or additional context.
 In this way, it can leverage off-the-shelf LLMs effectively without code-specific fine-tuning (Dong et al., 2025; Pan et al., 2025).
-However, to the best of our knowledge, there has been limited exploration of LLMs for code generation in the context of low-resource or DSLs such as those used in ZKPs.
+However, to the best of our knowledge, there has been limited exploration of LLMs for code generation for low-resource programming languages (LRPL) or DSLs especially for ZKP DSLs.
 Benchmark.
-As an important part of LLM-based code generation research, benchmarks play a crucial role in understanding model performance and guiding future improvements.
+Benchmarks are a crucial component of LLM-based code generation research, playing a vital role in understanding model performance and guiding improvements.
 Currently, their accuracy and reasoning ability are usually evaluated on standard benchmarks on mainstream programming languages such as HumanEval (Chen et al., 2021) and MBPP (Research, 2021).
-However, evaluation tools for low-resource and DSLs remain limited, hindering comprehensive assessments of LLM performance in these contexts (Joel et al., 2024; Giagnorio et al., 2025).
+However, evaluation tools for LRPLs and DSLs remain limited, hindering comprehensive assessments of LLM performance in these contexts (Joel et al., 2024; Giagnorio et al., 2025).
 Code generation for ZKP programs exemplifies the challenges and opportunities in this space.
-Though benchmarks for low-resource languages (Cassano et al., 2022; Yang et al., 2023) and DSLs (Liu et al., 2023; Abukhalaf et al., 2023) exist, to our knowledge, no standardized benchmark suite specifically targets ZKP code generation, requiring tailored benchmarks and evaluation strategies.
+Though benchmarks for LRPLs (Cassano et al., 2022; Yang et al., 2023) and DSLs (Liu et al., 2023; Abukhalaf et al., 2023) exist, to our knowledge, no standardized benchmark suite specifically targets ZKP code generation, requiring tailored benchmarks and evaluation strategies.
 3. Measuring LLM Capability for ZK Code Generation
-As discussed in Sec.
-1 and Sec.
-2.1, ZK programming follows a proof-oriented model: developers translate high-level specifications into algebraic constraints over a finite field and produce proofs of satisfiability.
-This shift yields the workflow depicted in Fig.
-1, from constraint specification to implementation by composing gadgets and finally proof generation.
-We therefore assess two capabilities essential for end-to-end success: ➀ language and toolchain knowledge and ➁ gadget-level competence.
+As discussed in § 1 and § 2.1, ZK programming follows a proof-oriented model: developers translate high-level intent into algebraic constraints over a finite field and produce proofs of satisfiability.
+This shift yields a unique programming workflow depicted in Fig.
+1, where developers construct the verification relation by wiring together algebraic primitives (mathematical gates that enforce logical invariants over finite fields).
+While witness computation still resembles conventional imperative programming, the developer’s primary responsibility shifts to the rigorous specification of these verification constraints.
+To evaluate LLM proficiency in this critical domain, we assess two foundational capabilities essential for end-to-end success: ➀ language and toolchain knowledge and ➁ algebraic primitive competence.
   1. (1)
     Language and Toolchain Knowledge.
-Developers must understand syntax, typing rules, semantics of field operations, and the structure of witness assignments.
-Weakness in any of these areas routinely leads to programs that cannot be compiled, or programs that compile but encode an invalid relation.
+Developers must master domain-specific syntax, strict typing rules, and the semantics of field operations to structure valid constraint specification and witness assignments.
+Weakness in any of these areas routinely leads to programs that cannot be compiled or produce an invalid relation that renders the proof meaningless.
   2. (2)
-    Gadget-level Competence.
-Many basic operations in ZKP programs are not natively attainable in the programming frameworks.
-Instead, developers must rely on a library of pre-defined gadgets that encapsulate these operations with optimized implementations.
-Gadget-level competence refers to the ability to effectively utilize these gadgets in the context of ZKP development and, when necessary, implement new ones from scratch.
-Taken together, these two capabilities form the foundation of end-to-end generation of ZK programs.
-Without reliable language and toolchain knowledge, LLMs are unable to produce compilable programs, while without gadget-level competence, the resulting programs cannot faithfully encode the intended relations.
-In practice, developers move between applying syntax and compilation rules to structure programs and assembling or extending gadgets to realize non-trivial constraints, and for LLMs, success also hinges on integrating these skills that bridge surface-level language fluency with the deeper ability to model arithmetic and logical constraints.
-Existing benchmarks, however, primarily target mainstream programming models, leaving this special integration of capabilities unexplored.
-This gap motivates our study: to systematically measure and evaluate LLM capability for ZK code generation.
-To study that, our evaluation pipeline introduces ZK-Eval, which explicitly measures both dimensions and assesses how they combine in end-to-end generation tasks.
+    Algebraic Primitive Competence.
+Algebraic primitives are the atomic building blocks of ZK constraint systems.
+Since arithmetic circuits lack native support for high-level operations, developers combine relational (e.g., comparators), logical (e.g., bitwise gates), and arithmetic (e.g., division, inversion) primitives to express logic.
+Competence in these atomic units is a prerequisite for constructing sound, higher-level application-specific constraint systems.
+These capabilities constitute two core dimensions of end-to-end ZK program generation.
+Without language and toolchain knowledge, LLMs cannot produce compilable programs; without algebraic primitive competence, programs cannot faithfully encode intended relations.
+Since ZK programming requires integrating language-level fluency with deeper constraint modeling, existing mainstream benchmarks (Chen et al., 2021; Jain et al., 2024; Research, 2021) leave this integration unexplored.
+We introduce ZK-Eval to systematically measure both dimensions and their integration in end-to-end tasks.
 3.1. ZK-Eval Benchmark
-In this section, we present the curation of ZK-Eval.
-Instead of focusing solely on end-to-end program generation, we aim to measure the individual competencies that contribute to successful ZK code generation.
-Below, we outline the key components of ZK-Eval.
-Figure 3. The Design of MCQ Benchmark Evaluating LLM’s Knowledge on ZK Languages.
+To realize this benchmark design, we structure ZK-Eval in three stages, each targeting one capability and their integration. We illustrate the design overview in Fig. 3.
+Figure 3. The Three-Stage Evaluation Design of the ZK-Eval Benchmark.
 3.1.1. Language and Toolchain Knowledge
-We first evaluate whether LLMs possess the foundational knowledge required to author programs in ZK languages.
-To construct this evaluation, we curate a knowledge base covering the two most widely used languages, Circom and Noir.
+We first evaluate whether LLMs possess the foundational knowledge required to author ZK programs.
 As shown in the left side of Fig.
-3, sources are collected through automated web crawling of official documentation, including language specifications, tutorials, and API references (e.g., circomlib (iden3, 2025) for Circom and the Noir standard library).
-In addition, we manually gather publicly available examples from GitHub repositories (TEZCAN, 2025b; Sirin, [n.
-d.]; Contributors, [n.
-d.]; Thor, [n.
-d.]) and online tutorials (TEZCAN, 2025a) to ensure coverage of idiomatic usage.
-Based on this knowledge base, as shown in the middle of Fig.
-3, we design a pipeline to automatically generate multiple-choice questions, a format widely adopted in prior evaluations of language models (Hendrycks et al., 2020; Myrzakhan et al., 2024; Zellers et al., 2019; Clark et al., 2018).
+3, sources are collected through web crawling of official documentation, including language specifications, tutorials, and API references (e.g., Circom’s circomlib (iden3, 2025) and the Noir’s stdlib).
+Additionally, we gather publicly available pedagogical examples from GitHub repositories (TEZCAN, 2025b; Sirin, ; Contributors, ; Thor, ) and online tutorials (TEZCAN, 2025a) to increase coverage of idiomatic usage.
+This yields 51 documents for Circom and 40 for Noir.
+Based on this knowledge base, we then design a pipeline to automatically generate multiple-choice questions, a format widely adopted in prior evaluations of language models (Hendrycks et al., 2020; Myrzakhan et al., 2024; Zellers et al., 2019; Clark et al., 2018).
 Inspired by earlier work on LLM-driven question generation (Liu et al., 2020; Abbasiantaeb et al., 2024; Long et al., 2024; Ji et al., 2025), our pipeline iteratively samples subsets of documents and provides them to a large model, which composes candidate questions and answers.
 This process is repeated until the full document collection has been exhaustively covered, producing a diverse set of drafts.
-All generated questions are then subjected to rigorous expert validation (right side of Fig.
-3).
-Human reviewers examine each question for factual correctness, clarity, and alignment with the intended level of difficulty.
-Cross-checking is performed among the authors to ensure agreement, and questions that do not meet these standards are either corrected or discarded.
-Topic coverage is also verified to ensure that all major aspects of the knowledge base are represented.
-We categorize the questions into four groups: basic syntax (language grammar such as datatypes, statements, operators, control flows, signals and buses), advanced topics (e.g., taggings, traits, generics), API references (standard library templates and functions), and compiler principles (translation from source code to arithmetic circuits).
-A summary of the questions and representative examples is shown in Fig.
-4.
-As shown, the final benchmark consists of 172 questions for Circom and 164 for Noir, with a balanced distribution across the four categories.
-The full set of questions is provided in the Artifact.
-Figure 4. Illustrative MCQ on Basic Circom Syntax with a Summary of Knowledge Categories.
-Figure 5. The Design of Gadget Benchmark on LLM’s Ability to Use or Implement Gadgets to Encode Constraints.
-3.1.2. Using and Implementing Gadgets to Encode Constraints
-The second dimension of ZK-Eval assesses the ability of LLMs to use and implement gadgets to encode constraints, which are the fundamental building blocks of ZK programs.
-To construct this evaluation, we curate a representative set of gadgets from three complementary sources, as shown in Fig.
-5.
-First, circomlib (iden3, 2025), the official standard library of Circom, provides canonical implementations of common gadgets that developers routinely instantiate.
-Second, zk-kit (Nico Serrano, [n.
-d.]) extends this idea to a broader ecosystem, offering more reusable gadget implementations across multiple ZK languages.
-Third, as an important observation, many gadgets share similar logical and arithmetic structures with primitives in SMT solvers.
-To capture this perspective, we leverage the Z3Py (Research, [n.
-d.]) interface to the Z3 SMT solver (De Moura and Bjørner, 2008).
-These sources span both language-specific and language-agnostic perspectives, ensuring that our benchmark captures the core constraint patterns encountered in practice.
-The resulting gadget set consists of 35 gadgets covering a wide range of categories: (i) basic checks, including range tests and comparisons; (ii) logical gates, such as AND, OR, XOR, and equality; (iii) arithmetic operators, including addition, subtraction, multiplication, division, exponentiation, modular reduction, and inversion; and (iv) composite operators, such as sum, distinct, and multiplexors (see the full list in the Supplementary Material).
-This collection is sufficiently expressive to simulate the primitive layer of typical ZK development workflows: any non-trivial circuit can be realized through the composition of these gadgets.
-For every gadget in the set, we author a concise task description.
-Test cases are crafted to validate both completeness: accepting input assignments and soundness: rejecting assignments.
-For logical gates, this corresponds to enumerating the truth table; for arithmetic and composite gadgets, it is manually constructed based on their semantics.
-These tests ensure that correctness is evaluated not only in terms of producing the right outputs, but also in enforcing the intended constraints.
-Figure 6. The Design of End-to-end ZK Program Generation Benchmark Adapted from HumanEval.
+All generated questions are then subjected to rigorous expert validation.
+Three human reviewers independently assess each item for factual correctness, clarity, and alignment with the intended difficulty, with disagreements resolved through discussion to consensus.
+Cross-checking by the authors ensures consistency across the set.
+Topic coverage is also verified to ensure that all major aspects (as listed in the official documents) of the studied language are represented.
+We categorize the questions into four groups: basic syntax (language grammar such as datatypes, statements, operators, control flows, signals and buses), advanced topics (e.g., taggings, traits, generics), API references (standard library APIs), and compiler principles (translation from source code to arithmetic circuits).
+As shown in Fig.
+3, the final benchmark consists of 172 questions for Circom and 164 for Noir, with a balanced distribution across the four categories.
+Representative examples are provided in the Supplementary Material and the full set of questions is provided in the Artifact.
+3.1.2. Implementing and Using Algebraic Primitives to Specify Constraints
+We assess the capability of LLMs to implement and utilize algebraic primitives as the building blocks of ZK constraint systems.
+Different from higher-level cryptographic (e.g., hash functions, signatures) and application-level (e.g., account-balance checks, order-matching validity) APIs available in external libraries, these encapsulate the core relational logic that developers must wire to define custom circuit behavior.
+Technical Difficulty.
+We clarify that, constructing a benchmark that reflects the fundamental algebraic patterns essential to real-world ZK programs is challenging.
+In existing codebases, core relational logic is often entangled with application-specific boilerplate, making it difficult to isolate and evaluate the underlying algebraic reasoning.
+To decouple and isolate these foundational patterns, we systematically curate a set of primitives from widely adopted libraries.
+This process, illustrated in Fig.
+3, draws from three complementary sources: circomlib (iden3, 2025), zk-kit (Nico Serrano, ), and Z3Py (Research, ).
+  1. (1)
+    Standardized Libraries. We first include canonical primitives from circomlib and zk-kit, which represent the production standard for routine circuit instantiation.
+  2. (2)
+    Agnostic Extraction.
+From these repositories, we manually extract and deduplicate algebraic primitives while filtering out application-specific composites.
+This ensures the benchmark captures the universal building blocks of ZK logic rather than application-specific patterns.
+  3. (3)
+    Aligned Completeness.
+To ensure comprehensiveness, we cross-reference our set with the Z3Py SMT-solver interface and incorporate primitives with similar structures, guaranteeing completeness for algebraic constraint specification.
+The resulting benchmark consists of 35 algebraic primitives covering a wide range of categories: (i) basic checks, including range tests and comparisons; (ii) logical gates, such as AND, OR, XOR, and equality; (iii) arithmetic operators, such as addition, multiplication, integer division, exponentiation, modular reduction, and inversion; and (iv) composite operators, such as sum and multiplexers (see the full list in the Supplementary Material).
+This collection serves as a complete basis for constraint specification; as these algebraic primitives represent the atomic operations of arithmetic circuits, any non-trivial verification relation can be realized through their rigorous composition.
+Practical Relevance.
+To check the practical relevance of our benchmark, we surveyed 37 production-grade ZK repositories comprising 408 source files.
+Our analysis confirms that the selected algebraic primitives are highly representative of real-world circuit logic: we achieve 97.30% and 94.59% coverage for core arithmetic and relational primitives, respectively, which are universal to production ZK systems.
+While logical (70.27%) and specialized composites (16.22%) appear less frequently, they remain mathematically essential for logical completeness and the construction of complex verification relations.
+Notably, it includes every one of the top 10 most frequently used primitives identified in the survey, establishing a representative foundation.
+Task Setup.
+For each primitive, we author a concise task description that asks the LLM to implement the primitive’s functional logic and use it to specify the intended algebraic relation.
+Following established practice on test-case curation (Chen et al., 2021; Jain et al., 2024), we craft test-cases that validate both completeness (accepting valid witnesses) and soundness (rejecting invalid assignments).
+For finite primitives (e.g., logical xor), test-cases exhaustively enumerate the entire truth table.
+For primitives with larger domains, we combine automated random input generation with edge-cases (e.g., field overflows, zero-value or out-of-domain inputs), yielding at least 100 cases per primitive.
+To validate test rigor, we injected artificial errors into ground-truth implementations (like a mutation testing process); our tests detected 87.04% of these faults, confirming their effectiveness at catching subtle mistakes.
 3.1.3. End-to-End ZK Program Generation
-The third dimension of ZK-Eval evaluates the capability of LLMs to perform full end-to-end ZK program generation, which requires the model to integrate language proficiency and gadget-level competence to assemble complete programs that encode non-trivial problem specifications.
-We build the benchmark on the top of the widely used HumanEval dataset (Chen et al., 2021) with ZKP-specific adaptions, as shown in Fig.
-6.
-In particular, we exclude tasks involving (i) string manipulation, (ii) floating-point arithmetic, and (iii) outputs of variable length.
-Strings and floating-point numbers are not natively supported by Circom or Noir, while variable-length outputs are incompatible with the static nature of arithmetic circuits.
-Following prior work (Myrzakhan et al., 2024), each task is automatically annotated by an LLM and verified by experts to ensure accurate filtering, yielding a set of 68 standalone tasks suitable for ZK program generation.
-To reflect the verification nature of ZK programs, we further reformulate the problem statements.
-Original HumanEval tasks are expressed as construction problems, in which a function computes an output $S=P(I)$ for input $I$.
-In contrast, ZK programs implement verification problems, where the goal is to check whether a candidate solution $S$ satisfies the intended relation for $I$.
-We therefore restate each task as a verification function $V(I,S)\in\{0,1\}$, which outputs one if and only if the candidate solution is correct.
-For each task, we construct an executable Python oracle that implements the verification semantics independently of any ZK framework, serving as a ground truth reference.
-Following prior work (Yuan et al., 2023; Siddiq et al., 2023), we then instruct an LLM to generate two Python scripts for test case generation: one producing input-solution pairs $(I,S)$ that should be accepted, and the other producing pairs that should be rejected.
-This dual design is crucial, as existing code-generation benchmarks provide only accepting examples (verifying acceptance of correct input-solution pairs) but do not evaluate rejection behavior.
-This ensures that generated programs are tested both for soundness (rejecting invalid ones) and completeness (accepting valid ones).
-To ensure the benchmark offers a rigorous and faithful assessment, all test cases are first executed against the Python oracle and then validated and cross-checked by authors for clarity, correctness, and corner-case coverage, following practices in prior benchmarks (Liu et al., 2025; Zhang et al., 2024b; Wang et al., 2022a).
-Whenever inconsistencies are detected at any stage, the corresponding step is re-executed or manually corrected.
-If corner-case is found lacking, the test case generator is revised accordingly.
+The final stage assesses LLMs to generate complete and correct ZK programs to verify computational statements.
+Building upon the first two stages, it evaluates the systemic integration of these skills to construct sound, multi-primitive constraint systems that faithfully encode high-level verification relations.
+Benchmark Design Rationale.
+To ensure a rigorous and trustworthy assessment in the absence of a standardized ZK-specific generation dataset, we leverage canonical algorithmic baselines from the well-established HumanEval dataset.
+We posit that ZK programming represents a fundamental paradigm shift from imperative computation to algebraic verification rather than a shift in the application domain itself.
+Since the core logic of production-grade ZK applications (e.g., private finance, decentralized identity, and state-transition proofs) relies on universal algorithmic patterns (e.g., sorting, membership checks, and conditional branching), adapting these general algorithmic tasks allows us to isolate the gap on computation-to-verification translation without the confounding variables of application-specific requirements (e.g., application-specific API calling).
+To reflect the unique semantics of ZK proofs, we systematically transform the original benchmark through three distinct adaptations, as illustrated in the right column of Fig. 3:
+  1. (1)
+    Compatibility Filtering.
+We exclude tasks involving string manipulation and floating-point arithmetic which are not supported by DSLs.
+Variable-length output tasks are also excluded since they’re incompatible with the static nature of arithmetic circuits.
+Their removal ensures that evaluation focuses on the model’s constraint-reasoning ability rather than DSL limitations.
+This yields a curated set of 68 standalone tasks suitable for ZK synthesis.
+  2. (2)
+    Verification Reformulation.
+We reformulate the natural language description of each computational task (defined as finding $S=P(I)$ for input $I$) into a verification problem.
+The LLM must generate a verification function $V(I,S)\in\{0,1\}$ that returns $1$ if and only if the candidate solution $S$ satisfies the intended relation for input $I$.
+  3. (3)
+    Oracle Adaptation.
+Correspondingly, the ground-truth oracle solutions are rewritten from computational functions ($S=P(I)$) to verification forms ($V(I,S)$).
+To ensure these serve as a reliable ground truth, we adopt a declarative implementation style; for instance, a sorting task is verified by checking permutation and monotonicity predicates rather than re-executing a sorting algorithm.
+To align with the semantics of ZKPs, these oracles are restricted to finite field arithmetic (mod $p$) and avoid non-deterministic execution paths.
+Each oracle was manually authored and cross-validated against the original computational benchmarks to ensure its reliability as a reference for test-case preparation.
+Quality Assurance.
+Following established practices in benchmark construction (Yuan et al., 2023; Siddiq et al., 2023; Liu et al., 2025; Zhang et al., 2024b; Wang et al., 2022), we implement a rigorous validation process to ensure the Soundness and Completeness of generated programs.
+While traditional datasets provide only “Acceptance Cases” (Completeness), ZK programs require “Rejection Cases” to ensure the circuit is properly constrained (Soundness).
+Thus we implement test-case generators to produce input-solution pairs that a correct verifier must reject, ensuring soundness validation.
+We cross-validate task descriptions, adapted oracles, and test-case generators through manual review and injection of edge cases (e.g., all-zeros, field-overflows, out-of-domain inputs, etc.).
+Collectively, this process yields 100 test cases per task with an average distribution of 54.41% acceptance and 45.59% rejection cases.
+To quantify benchmark quality, we injected syntactic mutations into the ground-truth oracles (mutation-testing); our test suites detected 86.42% mutations, demonstrating high sensitivity to subtle errors.
 3.2. Measurement Settings
 Using ZK-Eval, we measure LLMs’ competencies on ZK code generation. Our study is guided by the following three research questions (RQs):
-  • RQ1: Language & Toolchain Knowledge. Do LLMs understand the syntax, semantics, and standard libraries of Circom and Noir?
-  • RQ2: Gadget-Level Competence. Can LLMs reliably use or implement constraint gadgets, the building blocks of ZK programs?
-  • RQ3: End-to-End Generation. Can LLMs integrate these skills to produce complete, correct ZK programs?
+  • RQ1: Language & Toolchain Knowledge. Do LLMs understand the syntax, semantics, and standard library APIs of Circom and Noir?
+  • RQ2: Algebraic Primitive Competence. Can LLMs reliably implement and use algebraic primitives, the building blocks of ZK programs, to specify correct constraints?
+  • RQ3: End-to-End Generation. Can LLMs integrate these skills to produce complete, correct ZK programs on the intended computational relation?
 Each RQ corresponds directly to one component of ZK-Eval.
 For RQ1, models are tested on curated multiple-choice questions covering language knowledge.
-For RQ2, they are instructed to implement gadget tasks which are then automatically validated against accepting and rejecting test cases.
-For RQ3, they attempt end-to-end program generation from reformulated HumanEval tasks, with correctness judged by soundness and completeness against dual test suites.
+For RQ2, they are instructed to implement and use algebraic primitives which are then automatically validated against acceptance and rejection test cases.
+For RQ3, they attempt end-to-end program generation to realize verification constraints for general computational problems, whose correctness is judged by soundness and completeness against dual test suites.
 Model Selection.
 We evaluate 4 representative LLMs: OpenAI’s reasoning models GPT-o4-mini and GPT-o3, along with DeepSeek-V3 and Qwen3.
-This mix of reasoning and open-source non-reasoning ones allows us to assess both family-specific and general trends.
-All models are run under identical prompts and budgets, with 10 samples per task to estimate pass rates and reduce variance.
+This mix of reasoning and open-weight non-reasoning ones allows us to assess both family-specific and general trends.
+All models are run under identical prompts and budgets, with 10 samples per task to estimate Pass@1 rates and reduce variance.
 Figure 4. Accuracy of LLMs and Human Experts on the MCQ Benchmark for ZK Language Knowledge.
 Figure 5. LLM Performance on End-to-end ZK Benchmarks. Annotated with average accuracy.
-Figure 9. Average Error Distribution of Gadget Implementations Across Languages, Types, and Causes.
+Figure 6. Average Error Distribution of Primitive Implementations Across Languages, Types, and Causes.
 3.3. (RQ1) Language & Toolchain Knowledge
 Fig.
-3.2 summarizes performance on the multiple-choice question set.
+3.2 summarizes performance.
 Reasoning models perform best, with GPT-o4-mini reaching 88.1% and GPT-o3 87.2%.
-Open-source non-reasoning models perform noticeably worse, with DeepSeek V3 at 79.5% and Qwen3 at 78.9%.
+Open-weight non-reasoning models perform noticeably worse, with DeepSeek-V3 at 79.5% and Qwen3 at 78.9%.
 The human expert for reference scores 88.7%, similar to the reasoning models.
 Overall, LLMs demonstrate strong competence in ZK language knowledge, achieving accuracy within a few percentage points of the human expert.
-Finding ①: On Circom and Noir, LLMs display strong competence in ZK language knowledge. Reasoning models surpass open-source ones and perform similar to human experts.
+Finding ①: On Circom and Noir, LLMs display strong competence in ZK language knowledge. Reasoning models surpass open-weight ones and perform similar to human experts.
 We further examine performance across question categories and languages.
@@
 In contrast, accuracy drops on advanced topics such as traits, signal taggings, and generics, which extend expressiveness but are less commonly encountered.
-The lowest performance is observed on compiler principles, which probe understanding of how source programs are transformed into arithmetic circuits.
-These results indicate that while LLMs are proficient in surface-level language usage, they struggle with deeper aspects of language design and compilation.
+The lowest performance is observed on compiler principles, which probe understanding of how source programs are transformed into arithmetic circuits, indicating that they struggle with deeper aspects of arithmetic circuits and constraint semantics on finite fields.
 Notably, there is no significant difference between Circom and Noir, suggesting that model competence reflects general exposure to ZK languages rather than language-specific familiarity.
 Finding ②: LLMs excel at surface-level knowledge such as syntax and API references, but exhibit weaknesses in advanced language features and compiler principles.
-These findings are consistent with the nature of the knowledge base: official documents, standard libraries, public tutorials, and open-source examples are all widely accessible training material to LLMs.
-Therefore, it is anticipated that LLMs can achieve strong performance on such content.
-3.4. (RQ2) Gadget-Level Competence
-Building on the language knowledge observed in RQ1, we next assess whether LLMs can effectively apply this knowledge in ZK programming, specifically in using or implementing gadgets.
-Model outputs were analyzed with automated scripts that classified errors based on compiler diagnostics and test outcomes.
-From their distinct linguistic and functional effects, we grouped errors into six categories: (i) constraint enforcement, where models failed to encode relations using equality constructs ( === in Circom, assert in Noir); (ii) modules, involving incorrect instantiation or usage of library components; (iii) signal handling, specific to Circom, where mistakes occurred in declaring, assigning, or referencing signals; (iv) syntax & parsing, where the generated code violated grammar rules; (v) typing & resolution, where data types or variables were used inconsistently or incorrectly; and (vi) semantic errors, where code compiled but failed functional correctness tests.
+These findings are consistent with the nature of the knowledge base: official documents, public tutorials, and examples are all widely accessible training material to LLMs.
+Thus, it is anticipated that LLMs can achieve strong performance on such content.
+This suggests that (i) explicit augmentations on language grammars or official documentations may be unnecessary (ii) LLMs exhibit strong potential for ZK programming tasks but require further evaluations to reveal the bottlenecks.
+3.4. (RQ2) Algebraic Primitive Competence
+Language knowledge enables syntactic fluency, but ZK programming fundamentally requires the ability to specify algebraic constraints.
+We next assess whether LLMs can use or implement algebraic primitives, the atomic elements of constraint systems.
+Specifically, model outputs were analyzed with automated scripts that classified errors based on compiler diagnostics and test outcomes.
+From their distinct linguistic and functional effects, we grouped errors into six categories: (i) constraint enforcement, where models failed to encode relations using equality constructs ( === in Circom, assert in Noir); (ii) modules, involving incorrect instantiation or usage of library components; (iii) signal handling, specific to Circom, where mistakes occurred in declaring, assigning, or referencing signals; (iv) syntax & parsing, where the generated code violated grammar rules; (v) typing & resolution, where data types or variables were used inconsistently or incorrectly; and (vi) functional errors, where code compiled but failed functional correctness tests.
 Fig.
-9 shows the results on the gadget benchmark.
-Performance is consistently weak: even the best-performing category, logical operations, reached only 52% accuracy in Circom and 49% in Noir.
-Comparison gadgets scored lower (37% and 20%), while arithmetic and composite gadgets dropped to extremely low accuracy.
-This contrasts sharply with the strong results on language-level knowledge: when asked to operationalize this knowledge into gadgets, LLMs largely fail to produce correct, testable implementations.
-This gap reflects the difficulty of code generation for low-resource DSLs (Joel et al., 2024), exacerbated in ZK programming by limited and fragmented resources (Sheybani et al., 2025).
-Finding ③: On Circom and Noir languages, LLMs perform consistently weak on using or implementing constraint gadgets across all categories.
-Typing and resolution errors are the most frequent.
-Noir enforces strict type matching in binary operations, disallowing implicit casts between fields and integers, while Circom distinguishes sharply between variables and signals.
+6 shows the results and performance is consistently weak: even the best-performing category, logical operations, reached only 52% accuracy in Circom and 49% in Noir.
+Comparison scored lower (37% and 20%), while Arithmetic and Composites dropped to extremely low accuracy.
+This contrasts sharply with the strong results on language-level knowledge: when asked to operationalize this knowledge into specifying constraints with algebraic primitives, LLMs largely fail to produce testable, correct implementations.
+This gap reflects the difficulty of code generation for low-resource DSLs (Joel et al., 2024), exacerbated in ZK programming by the inherent programming paradigm shift (Sheybani et al., 2025).
+Finding ③: On Circom and Noir languages, LLMs perform consistently weak on implementing and using algebraic primitives to specify constraints across all categories.
+Typing & resolution errors are the most frequent.
+Noir enforces strict type matching, disallowing implicit casts between fields and integers, while Circom distinguishes between variables and signals.
 Models frequently mix types, reference undeclared identifiers, or redefine variables, causing compilation failures.
 Signal handling errors, unique to Circom, are the next most common, including declaring signals inside control-flow, reassigning them multiple times, or using them before initialization.
-Constraint enforcement errors also recur: instead of encoding equalities with === or assert, models generate invalid forms such as non-quadratic constraints in Circom or deprecated constructs in Noir.
+Constraint enforcement errors also recur: models generate invalid constraint forms such as non-quadratic constraints (e.g., x * y * z === 1) or use off-circuit assertions.
 These issues stem from ZK-specific semantics absent in conventional languages.
-Performance further varies by gadget type, reflecting library support and algebraic complexity.
+Performance further varies by primitive type, reflecting library support and algebraic complexity.
 Logical operations and comparisons, being boolean primitives or supported by libraries, are most tractable.
-Arithmetic and composite gadgets, such as exponentiation, modulo, floor division, or composites like sum, product, and distinct, lack native support and must be implemented from scratch.
+Arithmetic and Composites, such as exponentiation, modulo, floor division, or composites like sum and product, lack native support and must be implemented from scratch.
 These tasks demand reasoning over field arithmetic, careful handling of inverses and negatives, and explicit soundness constraints.
-The very low accuracy here underscores the inability of current models to generate gadgets beyond those directly provided in libraries.
-Finding ④: LLMs handle boolean and comparison gadgets better than arithmetic and composite ones, where ZK-specific semantics and limited library support demand non-trivial reasoning.
+The very low accuracy here underscores the inability of current models to generate implementations beyond those directly provided in libraries.
+Finding ④: LLMs handle boolean and comparison primitives better than arithmetic and composite ones, where ZK-specific semantics and limited library support demand non-trivial reasoning.
 3.5. (RQ3) End-to-End Generation
@@
 5 summarizes the results on the end-to-end generation benchmark.
-Compared to language knowledge and gadget-level tasks, performance drops substantially across all models.
+Compared to language knowledge and primitive-level tasks, performance drops substantially across all models.
 For Circom, accuracy is very low: GPT-o3 reaches only 20% semantically correct and 22% syntactically valid programs, while Qwen3 drops to 5% and 9%.
 Noir results are somewhat higher but still limited, with GPT-o4-mini at 32% and 40% and Qwen3 at 10% and 19%.
-Across all settings, semantic correctness lags behind syntactic validity, showing that many compilable programs still fail to encode the intended verification relation.
 Several trends emerge from these results.
 First, Noir tasks consistently yield higher success rates than Circom, likely due to Noir’s simpler type system and cleaner syntax, which reduce opportunities for errors such as signal misuse or type mismatches that are pervasive in Circom.
-Second, reasoning models substantially outperform open-source ones: GPT-o4-mini and GPT-o3 are 2–3 times more successful than DeepSeek V3 and Qwen3 across both languages, confirming the gap already observed in earlier tasks.
-Finally, the gap between syntactic and semantic correctness underscores that compilation alone is insufficient for assessing ZK program quality: many outputs are compilable but wrong.
-Finding ⑤: LLMs perform poorly on end-to-end code generation: reasoning models beats open-source ones; Noir is easier than Circom; many programs compile but remain incorrect.
+Second, reasoning models substantially outperform open-weight ones: GPT-o4-mini and GPT-o3 are 2–3 times more successful than DeepSeek-V3 and Qwen3 across both languages, confirming the model capability difference already observed in earlier tasks.
+Finally, the gap between syntactic and semantic correctness underscores that being compilable alone is insufficient for assessing ZK program quality: many outputs are compilable but suffer from under- or over-constrained relations.
+Finding ⑤: LLMs perform poorly on end-to-end code generation: reasoning models beats open-weight ones; Noir is easier than Circom; many programs compile but remain incorrect.
+This reveals the fundamental challenge of ZK programming: it requires not only language fluency and primitive competence, but also the ability to architect multi-primitive constraint systems that faithfully encode complex verification relations — a skill that current LLMs have not yet mastered.
 4. Augmenting LLMs for ZK Code Generation
-Our study reveals a gap between LLMs’ surface-level knowledge of ZK DSLs and their ability to construct correct gadgets or complete programs.
-Models can recall syntax and library references but struggle to translate this knowledge into valid constraint systems.
-To bridge this gap, we introduce ZK-Coder, an agentic framework for end-to-end ZK code generation.
+Based on the findings and gaps revealed from our empirical study in § 3, we introduce ZK-Coder, an agentic framework to augment end-to-end ZK code generation.
 We first summarize the design principles derived from our findings, then present the system.
 4.1. Design Principles Learned from the Findings
-Findings ① and ② show that LLMs possess a strong knowledge of ZK DSL syntax, semantics, and standard libraries, achieving near expert-level accuracy even on advanced features.
-This reflects their absorption of publicly available documentation and indicates their potential for ZK code generation.
-Despite this, limitations remain that motivate our further exploration.
-In Findings ③ and ④, we observe that this knowledge does not translate into reliable gadget-level generation.
-Generated code frequently suffers from typing errors, signal mismanagement, and faulty constraint enforcement, reflecting the scarcity of training examples with concrete gadget implementations.
-To bridge this gap between knowing the language and encoding correct programs, we introduce a sketch language in ZK-Coder that abstracts ZKP’s underlying constraint system.
-To further link them to curated hints, ZK-Coder leverages retrieval-augmented generation (RAG) (Gao et al., 2023) to guide gadget usage.
-Loosely speaking, RAG trades brittle memorization for grounding: once the needed gadgets are identified, retrieval supplies exact usage patterns that collapse the search space and curb gadget hallucinations.
-This strategy explicitly mirrors how developers consult a codebook and enables composition of verified snippets for rare gadgets that pretraining scarcely covers.
-Finally, Findings ③ and ⑤ highlight end-to-end program generation as the most challenging setting, where generated programs often fail due to syntactic errors or semantic mismatches.
+Findings ① and ② show that LLMs possess strong knowledge of ZK languages’ syntax, semantics, and standard library APIs, achieving near expert-level accuracy.
+Overall, we interpret that knowledge has been internalized during contemporary LLM’s pretraining stage, eliminating the need for grammar prompting or documentation access in our augmentation.
+However, important challenges remain that motivate our further exploration.
+In Findings ③ and ④, we observe that this knowledge does not translate into reliable primitive-level generation.
+Generated code frequently suffers from typing errors, signal mismanagement, and faulty constraint enforcement.
+Finding ⑤ also reveals these failures in end-to-end generation.
+To bridge this gap, ZK-Coder introduces a sketch layer that decouples abstract constraint modeling from low-level DSL implementation.
+By first formalizing relations in this intermediate representation, the framework can precisely map identified primitives to verified circuit templates and usage hints via retrieval-augmented generation (RAG) (Gao et al., 2023).
+Loosely speaking, this trades brittle memorization for grounding.
+Once the logical “what” is defined in the sketch, retrieval provides the exact “how” for DSL-specific wiring, collapsing the search space and enabling the reliable composition of even rare algebraic snippets.
+Finally, Findings ③ and ⑤ highlight end-to-end program generation as the most challenging setting, where generated programs often fail due to constraint-misspecification or under/over-constrained logic.
 These failures directly undermine the reliability and necessitate mechanisms to improve validity and faithfulness.
-To mitigate these failures, ZK-Coder introduces a correctness-centric interactive loop.
-The LLM proposes a program, queries a compiler as an oracle to check syntactic validity, and, once compilable, executes test cases as an oracle to check intended semantics.
-By observing diagnostics and counterexamples, ZK-Coder repairs its code and iterates the “generate-compile-test-repair” loop until both syntactic and semantic checks pass.
+To mitigate these failures, ZK-Coder introduces an interactive refinement loop.
+The LLM proposes a program, queries the compiler as an oracle to check constraint specification validity, and, once passes, self-supplies acceptance/rejection test cases to validate relation faithfulness.
+By observing diagnostics and counterexamples, ZK-Coder repairs its code and iterates the “generate-compile-test-repair” loop until these checks pass.
 This interaction transforms failures into actionable feedback and guides candidates toward valid, faithful programs.
@@
 Based on these design principles, we developed ZK-Coder, illustrated in Fig.
-10.
+7.
 ZK-Coder augments LLMs with a structured pipeline that bridges natural-language problem descriptions and executable ZK programs.
-It consists of three main stages: ➀ constraint formulation, ➁ constraint-guided analysis and retrieval, and ➂ interactive generation and repair.
-In ➀, problem statements are translated into lightweight sketch language.
-Then, in ➁, the sketch is analyzed to extract its constraints, which are used to retrieve relevant usage patterns on gadgets from a curated knowledge base.
-Finally, in ➂, the LLM generates the target ZK code using the retrieved hints, interactively repairing it with the help of syntax and semantic oracles until correctness is achieved.
-We detail the design of these stages in this section, and present an illustrative workflow with concrete code examples in the Supplementary Material.
+It consists of three main stages: ➀ constraint formulation, ➁ sketch-guided analysis and retrieval, and ➂ interactive generation and repair.
 Figure 7. Illustrative Overview on the Design of ZK-Coder.
@@
 Figure 8. The Abstract Grammar of the ZK Sketch Language.
-The first stage of ZK-Coder reasons and translates a natural-language problem description into a structured representation of ZKP programs.
-In accordance to our observations in Sec.
-3.4 as well as prior work on sketching and planning for code generation (Wang et al., 2024; Deng et al., 2024; Li et al., 2023; Jiang et al., 2024a; Wen et al., 2024a), we introduce a lightweight intermediate language, the ZK Sketch Language (ZKSL; Fig.
-11) that shares a similar essence with Python and Z3 ecosystem while abstracting away ZK-specific details.
-In ZKSL, a Program consists of a single verify function with typed parameters and a sequence of statements (Stmt); statements include assignments, assertions via engine.add($\Phi$), bounded loops and conditionals, and an optional return.
-Formulas $\Phi$ are built from relational atoms $E\bowtie E$ and gadget calls $G(\cdot)$.
-Its primitives (literals, arithmetic operators, etc.) align with the gadgets used in Sec.
-3.1.1.
-ZK-Coder prompts LLMs to generate ZKSL sketches from problem specifications.
-Then, to ensure well-formedness, ZK-Coder incorporates a checker that validates sketches on syntactic validity, type/arity consistency, and closure.
-If a sketch is invalid, feedback is provided to the LLM to refine its output.
-Since ZKSL is ZK-agnostic, it isolates constraint reasoning from ZK-specific syntax and semantics and reduces the generation complexity.
-4.2.2. Constraint-Guided Analysis and Retrieval.
-In addition to sketching and planning with ZKSL, RAG has been shown powerful to aid LLM-based code generation in low-resource, domain-specific languages (Joel et al., 2024).
-After obtaining a ZK-agnostic sketch, ZK-Coder analyzes its constraints and retrieves ZK-specific implementation guidance, bridging the abstract ZKSL specification and the concrete gadget implementations required in particular ZK DSLs.
-We traverse the ZKSL sketch, collecting each constraint introduced via engine.add($\cdot$).
-Operators are canonicalized: aliases are normalized (e.g., >= is normalized to GreaterEqThan) and operator variants are distinguished by arity and types (e.g., XOr: Bool $\times$ Bool $\rightarrow$ Bool vs.
-XOr: Field $\times$ Field $\rightarrow$ Field).
-Non-constraint constructs (e.g., folded constants, loop statements or conditional guards that emit no constraints) are discarded.
-The result is an ordered, typed set of constraint primitives corresponding to the required gadgets.
-To build a knowledge base for retrieval, we curate a collection of gadget and their usage examples either from the standard libraries of ZK DSLs or from verified implementations.
-With the knowledge base, ZK-Coder performs constraint-guided retrieval to fetch relevant usage patterns on gadgets for each constraint primitive.
-Retrieval enforces exact operator/type/arity matching, and hints are emitted in the ZKSL order to preserve program flow during generation.
-In this way, ZK-Coder effectively grounds generation in precise constraint analysis and structured retrieval.
-4.2.3. Interactive Generation and Repair.
-With the sketch generated in ➀ and the hints retrieved in ➁, ZK-Coder proceeds to the final stage of end-to-end code generation.
-As pointed out in prior work (Joel et al., 2024; Wang et al., 2022b; Bi et al., 2024; Skreta et al., 2023), introducing a feedback loop that leverages compilers and test executors as oracles can substantially improve code generation reliability.
-Motivated by this, ZK-Coder implements an interactive generate-repair loop to ensure both syntactic validity and semantic correctness of the generated ZK programs.
-The loop alternates between code generation, compilation checks, and execution-based testing, iteratively refining the program until it meets all correctness criteria.
-Given the ZKSL sketch specification and constraint-specific hints, the LLM produces a candidate program.
-Then, ZK-Coder submits the program to the ZKP compiler (i.e., Circom or Noir) for compilation.
-If ZK-Coder observes compilation errors (e.g., syntax errors, type mismatches, or illegal signal operations), it extracts the diagnostics and feeds them back to the LLM for repair.
-The LLM refines the program based on this feedback, producing a new candidate that is again submitted for compilation.
-This process repeats until a compilable program is obtained.
-With a compilable program, ZK-Coder generates a set of test cases and validates the program on these cases on the fly.
-If any test case fails, ZK-Coder collects the counterexample and feeds it back to the LLM for further repair.
-This loop continues until the program successfully passes all test cases or the retry limit is reached.
-This interactive approach in ZK-Coder converts failures into constructive feedback and augments the LLM’s ability to produce reliable ZK programs.
+The first stage of ZK-Coder translates natural-language requirements into a structured specification of verification constraints.
+Drawing on the success of planning-based code generation (Wang et al., 2024; Deng et al., 2024; Li et al., 2023; Jiang et al., 2024b; Wen et al., 2024b), we introduce the ZK Sketch Language (ZKSL), a Python-embedded Domain Specific Language (eDSL) designed to bridge the gap between high-level intent and low-level circuit implementation.
+ZKSL is architected to allow LLMs to formalize constraint logic using familiar Pythonic syntax.
+This isolates the mathematical “what” from the DSL-specific “how” by requiring the explicit modeling of constraints over typed inputs to mirror the verification nature of ZK programs.
+As formally defined [^1] We also provide concrete running examples in the Supplementary Material.
+in Fig.
+8, a ZKSL Program is structured around a verify function that accepts these typed inputs and executes a sequence of constraint-declaration statements via engine.add($\Phi$).
+These formulas are composed of relational atoms that correspond to the algebraic primitives identified in § 3.1.2.
+Crucially, all operations within ZKSL are defined over finite fields (modulo $p$) to faithfully reflect the mathematical foundation of ZK systems.
+This enables LLMs to express complex verification relations using standardized constructs, serving as a direct logical precursor to the final target code.
+It allows LLMs to decouple algebraic reasoning from the implementation-level details (e.g., signal wiring and constraint encoding) of specific target DSLs.
+ZK-Coder prompts the LLM to generate these constraint specification sketches directly from problem descriptions.
+To ensure the integrity of the sketch, the framework performs automated syntactic and type/arity consistency checks.
+If a sketch is found to be invalid, the error trace is fed back to the model for iterative refinement.
+By serving as a logical representation, ZKSL decouples the complex task of algebraic constraint reasoning from the low-level specifics of ZK DSLs (e.g., signal wiring and constraint specification formats), reducing the overall generation complexity.
+4.2.2. Sketch-guided Analysis and Retrieval.
+RAG is a powerful technique for grounding LLM generation in low-resource or domain-specific contexts (Joel et al., 2024).
+ZK-Coder utilizes the ZKSL sketch as a structural query to bridge the gap between abstract constraint specifications and code implementations.
+Since a ZK program is defined by a set of verification constraints composed of atomic algebraic primitives, the framework traverses the sketch to extract each relation declared via engine.add($\cdot$).
+To ensure precise retrieval, operators are canonicalized by normalizing aliases (e.g., >= becomes GreaterEqThan) and distinguishing variants based on their arity and field-typing (e.g., differentiating XOr for Booleans vs.
+Fields).
+This stage isolates the constituent units within each constraint to determine the precise implementation hints required by the target DSL.
+To support this stage, we curate a knowledge base of verified implementations and idiomatic usage patterns for the algebraic primitives defined in § 3.1.2.
+Instead of text-similarity ranking, ZK-Coder then performs exact-match retrieval for each identified primitive.
+These patterns are emitted in the original ZKSL specification order to preserve the program’s logical flow during final code generation.
+This structured retrieval effectively grounds the generation process, collapsing the search space and significantly curbing the model’s tendency to hallucinate incorrect constraint semantics or signal-handling logic.
+4.2.3. Interactive Refinement Loop.
+While the sketch provides a logical blueprint, our findings in § 3.4 and § 3.5 reveal a persistent gap: many generated programs successfully satisfy DSL grammar rules but fail to faithfully encode the intended verification relation.
+Unlike mainstream software where errors often manifest as runtime crashes, ZK programs frequently yield under-constrained circuits or over-constrained systems, which silently accept invalid witnesses or sometimes reject valid ones.
+To bridge this gap, we implement an Interactive Refinement Loop as follows.
+  1. (1)
+    Constraint Construction Refinement.
+The candidate program is first submitted to the target compiler (e.g., Circom or Noir) to address DSL-specific failures identified in our study, such as signal mismanagement, illegal signal declarations in control flow, or constraint-encoding violations.
+We extract these diagnostic traces to iteratively refine the constraint wirings until they’re valid or reach the retry iteration threshold $N_{1}$ ($N_{1}$ is set as 8 in our experiments).
+  2. (2)
+    Relation Faithfulness Refinement.
+Once compilable, the framework validates constraint faithfulness.
+The agent generates representative “Acceptance” and “Rejection” test cases to verify that the circuit correctly enforces the intended relation, while evaluation test cases remain hidden to prevent data-leakage.
+If any test case fails, the counterexample is fed back to re-examine the algebraic constraints.
+This forms another refinement loop, and we set the retry threshold $N_{2}$ as 3 in our experiments.
+The above processes repeat until the program passes all functional checks or the budgets are exhausted, practically converting failures into guided refinements without manual intervention.
 5. Evaluation
@@
 Our evaluation is structured around the following research questions, with a focus on end-to-end generation.
-To clarify, we do not re-evaluate language knowledge (RQ1) or gadget-level coding (RQ2): the former is established in Sec.
-3.3, and the latter is handled by our retrieval mechanism in Sec.
-4.2.2.
-Given that said, we include an ablation study RQ5.
+To clarify, we do not re-evaluate language knowledge (RQ1) or primitive-level coding (RQ2): the former is established in § 3.3, and the latter is handled by our sketching and retrieval mechanism in § 4.
+Nevertheless, we include an ablation study (RQ5) to quantify the contribution of each component.
   • RQ4 (End-to-End Performance): How does ZK-Coder perform compared to baseline methods in generating correct ZK programs?
   • RQ5 (Ablation Study): What is the effect of each design component, including the ZK sketch language, constraint-guided retrieval, and the interactive repair loop?
-  • RQ6 (Robust Generalization): Beyond the standard benchmark, adapted HumanEval, how well does ZK-Coder transfer to newer and more challenging tasks?
-  • RQ7 (Failure Analysis): When ZK-Coder fails, what are the common failure patterns, and how might they inform future research?
+  • RQ6 (Robust Generalization): How well does ZK-Coder generalize to novel, complex computational tasks and real-world production-grade ZKP implementations?
+  • RQ7 (Failure Analysis): When ZK-Coder fails, what are common failure patterns, and how might they inform future research?
 5.1. Experimental Setup
 Our experiments are organized according to the research questions outlined above.
-For RQ4 (End-to-End Performance), we evaluate on the adapted HumanEval benchmark from ZK-Eval, using the two strongest OpenAI’s reasoning models from our measurement study: GPT-o4-mini and GPT-o3.
+For RQ4 (End-to-End Performance), we evaluate on the adapted HumanEval benchmark from ZK-Eval, using the same four models from our measurement study.
 This comparison establishes how well ZK-Coder outperforms baseline methods in generating correct and provable ZK programs.
-For RQ5 (Ablation Study), we conduct an ablation study on GPT-o4-mini, selectively disabling each design component of ZK-Coder, including the sketch language, constraint-guided retrieval, and the interactive repair loop.
+For RQ5 (Ablation Study), we conduct an ablation study on all four models, selectively disabling each design component of ZK-Coder, including the sketch language, constraint-guided retrieval, and the interactive repair loop.
 This aims to quantify their individual impact on overall performance.
@@
 This analysis highlights limitations of the current pipeline and suggests avenues for future research.
-All experiments are conducted under a fixed repair budget of up to 8 iterations for syntax checking and repair, and 1 iteration for semantic checking and comparison.
-Each task is executed 10 times to compute Pass@1 accuracy, ensuring statistical stability.
-The evaluation is conducted on a Linux (Ubuntu 22.04 LTS) server equipped with 256 GiB of RAM.
+All experiments are conducted under a fixed repair budget (see § 4.2.3). The evaluation is conducted on a Linux (Ubuntu 22.04 LTS) server equipped with 256 GiB of RAM.
 5.2. RQ4: End-to-End Performance
-We evaluate the end-to-end performance of ZK-Coder.
-To the best of our knowledge, since no prior work directly addresses zero-knowledge program generation from natural language, we establish a simple but meaningful baseline: directly prompting the model to produce circuit code from the task description and input specification, without any sketching, retrieval, or repair.
-This setting captures the raw capability of LLMs and provides a lower bound against which the contributions of ZK-Coder can be measured.
-During the evaluation, we meticulously document the performance for the pipeline in three key stages: (1) Sketch correctness measures whether the initial ZK sketch produced from the natural language description is correct by passing all hidden test cases; (2) Repair pass rate is conditioned on all correctly generated sketches, and reports the fraction of these that can be successfully repaired by the interactive repair loop; (3) Program correctness is conditioned on all successfully repaired programs, and measures the proportion that pass all hidden test cases.
+We evaluate ZK-Coder’s end-to-end performance.
+To the best of our knowledge, since no prior work directly addresses ZK program generation from natural language, we establish a simple but meaningful baseline: prompting the model to generate ZK code directly from task descriptions, with a summary of the target DSL’s grammar and few-shot examples.
+This setting captures the raw capability of LLMs as a lower bound.
+We provide more baselines in § 5.3.
+We meticulously document the performance for the pipeline in three key stages: (1) Sketch correctness measures whether the initial constraint sketch (in ZKSL) produced from the natural language description is correct by passing all hidden test cases; (2) Repair pass rate is conditioned on all correctly generated sketches, and reports the fraction of these that can be successfully repaired by the interactive repair loop; (3) Program correctness is conditioned on all successfully repaired programs, and measures the proportion that pass all hidden test cases.
 The final overall accuracy aggregates across the full pipeline, reflecting end-to-end program generation accuracy.
@@
 Correctness Pass Rate Correctness (Pass@1) Token (Pass@1) Token
-Circom GPT-o4-mini 94.12% 91.72% 96.59% 83.38% 2350.31 17.35% 436.98
-GPT-o3 97.30% 94.73% 96.82% 89.23% 2253.29 20.29% 433.35
-Noir GPT-o4-mini 94.12% 97.97% 97.66% 90.05% 648.95 32.21% 152.18
-GPT-o3 97.21% 97.43% 99.38% 94.12% 765.46 27.94% 195.34
-Table 3 summarizes the comparative performance of ZK-Coder and the baseline across both Circom and Noir.
-Across all settings, ZK-Coder achieves substantial gains over the baseline.
-On Circom, GPT-o4-mini and GPT-o3 reach overall pass rates of 83.38% and 89.23%, respectively, compared to only 17.35% and 20.29% for the baseline.
-Noir shows a similar trend, with GPT-o4-mini reaching 90.05% and GPT-o3 achieving the highest overall rate of 94.12%, far outperforming the corresponding baseline results of 32.21% and 27.94%.
+Circom GPT-o4-mini 94.12% 91.72% 96.25% 83.38% 2350.31 17.35% 436.98
+GPT-o3 97.30% 94.41% 95.64% 87.85% 2155.39 20.29% 433.35
+Deepseek V3 72.15% 92.77% 74.10% 49.60% 2958.72 7.35% 325.28
+Qwen 3 68.12% 79.07% 78.78% 42.43% 4979.72 4.90% 162.83
+Noir GPT-o4-mini 94.12% 97.97% 96.38% 88.87% 648.95 32.21% 152.18
+GPT-o3 97.94% 99.85% 100.00% 97.79% 750.63 28.38% 191.72
+Deepseek V3 74.88% 93.36% 94.31% 68.66% 628.69 16.18% 100.91
+Qwen 3 66.84% 91.88% 89.14% 54.74% 1260.43 11.27% 104.24
+Table 2 summarizes the comparative performance of ZK-Coder.
+Across all tested models and languages, ZK-Coder achieves substantial gains over the baseline.
+On Circom, GPT-o3 leads with 87.85% accuracy, while open-weight models (DeepSeek V3, Qwen 3) improve from below 10% baseline to 49.60% and 42.43%, respectively.
+On Noir, success rates are higher overall, likely due to its Rust-like syntax being better aligned with pretraining data.
+GPT-o3 still leads with 97.79% accuracy, and GPT-o4-mini achieves 88.87%.
+Open-weight models also show strong gains: Qwen 3 improves from 11.27% to 54.74%, and DeepSeek V3 from 16.18% to 68.66%.
 The per-stage breakdown highlights the robustness of the pipeline.
-Sketch correctness exceeds 94% across all configurations, showing that initial translation into ZK sketches is highly reliable.
-The repair loop maintains strong success rates, above 91% in Circom and above 97% in Noir, ensuring that most sketches can be generated into valid programs.
-Final program correctness surpasses 96% in all cases, confirming that the integration of sketching, repair, and semantic validation yields consistently executable and provable programs.
+OpenAI models show exceptional “zero-shot” capabilities in sketching constraints, while open-weight models exhibit lower correctness rates, suggesting that the leap from natural language to abstract constraint logic remains a challenge for these models.
+The repair loop maintains strong success rates, exceeding 90% in most settings, though Qwen 3 on Circom achieves 79.07%, indicating some difficulty in recovering from initial errors.
+Final program correctness is high across most configurations, though DeepSeek V3 and Qwen 3 show comparatively lower rates on Circom, suggesting challenges in implementing correct constraint logic.
+Overall, the integration of sketching, repair, and semantic validation yields consistently executable and provable programs across the majority of settings.
 We also note that ZK-Coder incurs a higher token cost compared to the baseline, which is expected since our agentic pipeline includes sketching, RAG and interactive repair.
-However, the cost remains modest in practice, with each task requiring fewer than 3,000 tokens on average, corresponding to less than 0.1 USD per task, making the gains in correctness highly cost-effective.
+However, the cost remains modest in practice, with each task requiring fewer than 5,000 tokens on average, corresponding to less than 0.1 USD per task, making the gains in correctness highly cost-effective.
 Answer to RQ4: ZK-Coder delivers a significant improvement in end-to-end performance over the baseline method.
@@
 5.3. RQ5: Ablation Study
-Figure 12. Accuracy Comparison of ZK-Coder Ablation Variants.
+Figure 9. Accuracy Comparison of ZK-Coder Ablation Variants (Averaged Across Models).
 To assess the role of each major component in ZK-Coder, we conduct an ablation study where individual modules are disabled or replaced. The following variants are considered:
-  • No RAG (V1): Disables constraint-guided retrieval augmented generation, forcing the model to rely solely on its internal knowledge.
-  • No Sketch (V2): Removes the intermediate ZK sketch reasoning and directly prompts the model to output full programs with RAG and the interactive repair loop.
-  • No Syntax Repair (V3): Eliminates the interactive repair loop for syntax errors.
-  • No Semantic Repair (V4): Retains syntax repair but disables semantic repair.
-  • Similarity-based RAG (V5): Replaces our constraint-guided retrieval with a naive similarity-based retrieval using surface-level embedding similarity.
-  • Only Repair (V6): Omits both retrieval and sketching, relying exclusively on the repair loop applied to direct generations.
+  • No RAG (V1): Disables sketch-guided RAG for algebraic primitives, forcing the model to rely solely on its internal knowledge.
+  • No Sketch (V2): Removes the intermediate ZK constraint sketching and directly prompts the model to output full programs with RAG and the interactive repair loop.
+  • No Compile-Repair (V3): Eliminates the interactive repair loop for compiler-identified errors.
+  • No Execute-Repair (V4): Retains compiler-guided repair but disables execution and repair.
+  • Similarity-based RAG (V5): Replaces our sketch-guided retrieval with a naive text-embedding similarity retrieval on the task description.
+  • Only Repair (V6): Omits both retrieval and sketching, relying exclusively on test-and-repair.
 Figure 9 presents the Pass@1 accuracy for each ablation variant across Circom and Noir.
 The results show that every component plays a critical role in the overall success of ZK-Coder.
-Removing RAG (V1) reduces accuracy from 83.38% to 67.21% on Circom, and from 90.05% to 89.22% on Noir.
-Disabling the sketch reasoning stage (V2) causes further degradation (71.32% on Circom, 87.50% on Noir), highlighting the value of sketching as a scaffold for task understanding and constraint reasoning.
-The most severe drop arises from removing syntax repair (V3), with accuracy falling to 35.00% on Circom and 41.39% on Noir, demonstrating that interactive repair is indispensable.
-Disabling semantic repair (V4) also leads to measurable reductions (81.18% on Circom, 87.99% on Noir), confirming that semantic validation contributes to robustness.
-Substituting constraint-guided retrieval with a naive similarity-based strategy (V5) yields weaker performance (69.87% on Circom, 89.41% on Noir), suggesting that constraint-guided retrieval is more effective than surface-level matching.
-Finally, the “Only Repair” variant (V6) performs poorly, especially on Circom (59.12%), indicating that repair alone cannot compensate for the absence of sketching and retrieval.
-Answer to RQ5: The ablation study shows that the strong performance of ZK-Coder relies on the synergy of sketching, constraint-guided RAG, and interactive repair.
-Removing or weakening any of these components causes substantial accuracy drops.
+Starting from a full-system average of 65.66% on Circom and 76.87% on Noir, we observe that removing RAG on primitives (V1) causes a significant drop to 53.04% and 72.00% respectively.
+Disabling the constraint sketching stage (V2) causes further degradation (46.37% on Circom, 64.63% on Noir), highlighting the value of sketching as a scaffold for task understanding and constraint modeling.
+The most severe drop arises from removing compile-repair (V3), with accuracy falling to 26.04% on Circom and 27.12% on Noir, demonstrating that interactive repair is indispensable.
+Disabling execute-repair (V4) also leads to measurable reductions (64.48% on Circom, 76.00% on Noir), confirming that semantic validation contributes to robustness.
+Substituting primitive-in-sketch-guided retrieval with a text-similarity-based strategy (V5) yields weaker performance (51.42% on Circom, 74.59% on Noir), suggesting that sketch-guided retrieval is more effective particularly for Circom.
+Finally, “Only Repair” (V6) achieves only 43.31% on Circom and 59.80% on Noir, showing that iterative refinement alone cannot overcome the “computation-to-verification” gap.
+Compared with V6, the full system achieves an up to 22.35% performance leap with modest token cost (20.46%-29.39%), confirming that constraint sketching and template retrieval outperform blind repair cycles.
+Answer to RQ5: The ablation study shows that the strong performance of ZK-Coder relies on the synergy of sketching, sketch-guided RAG, and interactive repair.
+Removing or weakening any of these components causes noticeable accuracy drops.
 5.4. RQ6: Robust Generalization
-To assess whether ZK-Coder sustains strong performance on harder, more recent, and contamination-free tasks, we evaluate on LiveCodeBench (Jain et al., 2024), a recently introduced benchmark designed to be contamination free.
-We select all tasks labeled as medium difficulty, filter and adapt them using the methodology introduced in Sec.
-3.1.3, resulting in a set of 34 tasks (details refer to the Supplementary Material).
-This offers a more challenging and realistic testbed for ZK program generation.
-Table 4. ZK-Coder Code Generation Performance on the Adapted LiveCodeBench Benchmark.
-Language ZK-Coder Baseline
-Circom 72.25% 2.97%
-Language ZK-Coder Baseline
-Noir 82.32% 14.72%
+Our adapted HumanEval serves as a controlled benchmark to evaluate the core computation-to-verification shift.
+To evaluate the robustness of ZK-Coder across varying levels of difficulty and practical contexts, we extend our analysis to two additional datasets: (i) LiveCodeBench (Jain et al., 2024), which focuses on high-complexity algorithmic challenges to assess the limits of the framework’s reasoning; and (ii) a set of Production Case Studies curated from 37 open-source ZKP repositories.
+To ensure the case studies are representative and avoid selection bias, we employed a reproducible filtering process targeting high-frequency circuit patterns such as Merkle verification, state nullification, etc.
+Details are provided in the Supplementary Material.
+Table 3. Generality Study on ZK-Coder’s Code Generation Performance (Averaged Across Models).
+Language LiveCodeBench Production Coding Pattern Case Studies
+ZK-Coder Repairing Baseline Simple Baseline ZK-Coder Repairing Baseline Simple Baseline
+Circom 44.10% 14.07% 1.72% 90.83% 83.33% 27.50%
+Noir 56.93% 33.95% 9.89% 92.09% 89.17% 31.67%
+We establish two baselines: the Simple Baseline (direct generation) and the Repairing Baseline (adding the interactive repair loop).
 Table 3 presents the results.
-ZK-Coder achieves 72.25% accuracy on Circom and 82.32% on Noir, dramatically outperforming the baseline scores of 2.97% and 14.72%, respectively.
-These results indicate that even on a harder and contamination-free benchmark, ZK-Coder continues to deliver strong accuracy and robust generalization.
-Answer to RQ6: ZK-Coder’s performance on LiveCodeBench confirms that the pipeline generalizes effectively to harder and contamination-free tasks.
+On LiveCodeBench, ZK-Coder achieves a dramatic performance leap, reaching 44.10% on Circom and 56.93% on Noir, significantly outperforming the Repairing Baseline (14.07% and 33.95%).
+On Production Case Studies, ZK-Coder reaches near-perfect success rates (90.83% on Circom and 92.09% on Noir).
+While the Repairing Baseline also performs strongly, this high baseline is expected, as these studies focus on standardized, commonly used verification patterns in a controlled context where the high-level logic is well-defined.
+Nevertheless, ZK-Coder achieves a noticeable gain by bridging the “last mile” of correctness, reducing subtle field-level constraint errors that persist in the baselines.
+Answer to RQ6: Results from LiveCodeBench and production case studies confirm that ZK-Coder generalizes effectively across a spectrum of challenges, from high-complexity verification logic to standardized production coding patterns.
 5.5. RQ7: Failure Analysis
-Table 5. Distribution of ZK-Coder’s Code Generation Failures.
-Language Benchmark Repair Budget Sketch False False Mixed False
-Exceed Incorrect Accept Reject Accept & Reject
-Circom Humaneval 46.09% 34.78% 0.87% 14.78% 3.48%
-LiveCodeBench 36.26% 37.36% 9.89% 15.38% 1.10%
-Noir Humaneval 17.33% 53.33% 5.33% 17.33% 6.67%
-LiveCodeBench 23.53% 58.82% 0.00% 17.65% 0.00%
-Table 5 reports the distribution of ZK-Coder’s failure cases across Circom and Noir on both HumanEval and LiveCodeBench.
+Table 4. Distribution of ZK-Coder’s Code Generation Failures (Averaged Across Models).
+Language ➀ Repair ➁ Sketched ➂ False Acceptance ➃ False Rejection ➄ Mixed False
+Budget Exceed Constraint Incorrect (Under-constrained) (Over-constrained) Acc / Rej
+Circom 31.67% 41.56% 1.08% 19.78% 5.92%
+Noir 12.60% 73.07% 1.67% 10.89% 1.78%
+Table 4 reports the distribution of ZK-Coder’s failure cases.
 Each column corresponds to a distinct failure mode.
 Repair Budget Exceed indicates cases where the repair budget was exhausted without producing any compilable program.
-Sketch Incorrect denotes that the generated ZKSL sketch failed hidden test cases, suggesting a misunderstanding of the task requirements.
-False Accept corresponds to programs that erroneously accept invalid witnesses, while False Reject refers to programs that incorrectly reject valid witnesses.
-Mixed False Accept & Reject captures programs that simultaneously exhibit both types of errors, reflecting severe constraint mis-specification.
-The data reveal that the two dominant sources of failure are Repair Budget Exceed and Sketch Incorrect.
-In both Circom and Noir, these two together account for over 80% of failures on both benchmarks.
-Repair Budget Exceed often arises when the model is unable to generate any compilable program, which due to gaps in pretraining exposure to ZK-specific constructs.
-Sketch Incorrect, on the other hand, point to deeper misunderstandings of the natural language description, where the task is not properly captured in the ZKSL representation.
-The less frequent categories (false accept, false reject, and mixed errors) typically stem from incorrect constraint enforcement when translating a valid sketch into a concrete circuit.
+Sketch Incorrect denotes that the generated constraint sketch failed hidden test cases, suggesting an incorrect specification of the verification constraints.
+False Acceptance corresponds to programs that erroneously accept invalid witnesses, while False Rejection refers to programs that incorrectly reject valid witnesses.
+Mixed False Acc / Rej captures programs that simultaneously exhibit both types of errors, reflecting severe constraint mis-specification.
+The data reveal that the two dominant sources of failures are ➀ and ➁.
+In both Circom and Noir, these two together account for over 70% of failures.
+Failure ➀ often arises when the model is unable to generate any compilable program, which is due to gaps in pretraining exposure to ZK-specific constructs.
+Failure ➁, on the other hand, reflects errors in correctly modeling the verification constraints from natural language descriptions.
+The less frequent categories (➂, ➃, ➄) typically stem from incorrect constraint enforcement when translating a valid sketch into a concrete circuit.
 These errors indicate local implementation mistakes rather than systemic failures of the pipeline.
-Answer to RQ7: In summary, Repair Budget Exceed and Sketch Incorrect constitute the most critical bottlenecks for robust end-to-end generation.
-The former highlights the need for more ZK-specific training data to improve raw generation capabilities, while the latter underscores the difficulty of grounding natural language descriptions in precise constraints.
+Answer to RQ7: In summary, Repair Budget Exceed and Sketched Constraint Incorrect constitute the most critical bottlenecks.
+The former highlights the need for more ZK-specific training data to improve raw generation capabilities, while the latter underscores the difficulty of grounding natural language descriptions in precise verification constraints.
 6. Related Work
 LLM for DSL Code Generation.
-Emerging research has investigated the use of LLMs for generating code in DSLs, a task made challenging by limited training data and specialized syntax (Joel et al., 2024).
-To mitigate these difficulties, grammar prompting has been proposed to guide LLMs with partial grammars and thereby improve the well-formedness of generated programs (Wang et al., 2023).
-RAG techniques further enhance DSL generation by supplying relevant API documentation or code samples (Pimparkhede et al., 2024; Bassamzadeh and Methani, 2024).
-For DSLs where correctness guarantees are critical, approaches based on formal verification (Councilman et al., 2025) and verified lifting (Bhatia et al., 2024) strengthen reliability by coupling generated code with machine-checkable proofs.
-Our work advances this line of research by focusing on DSLs for zero-knowledge proofs.
+Emerging research has investigated the use of LLMs for generating code in DSLs, a task made challenging by limited training data and specialized syntax (Joel et al., 2024; Giagnorio et al., 2025).
+To mitigate these difficulties, grammar prompting (Wang et al., 2023), RAG techniques (Pimparkhede et al., 2024; Bassamzadeh and Methani, 2024), iterative refinement (Ding et al., 2024), and knowledge transfer methods (Cassano et al., 2024; Mora et al., 2024) have been proposed.
+While most of these methods (except knowledge transfer methods, which target mainstream computational languages) can be adapted for ZKP code generation, our ablation study (§ 5.3) shows they remain ineffective.
+ZK-Coder advances this area by designing a system tailored to the constraint paradigm of ZKPs.
 Code Generation Benchmarks.
@@
 Software Engineering for ZKP
-The software engineering community has recently begun to explore tools and methodologies for ZKP development.
-Numerous researches have proposed to improve the security of ZKPs.
-Static analysis tools such as Circomspect (Dahlgren, 2022), ZKAP (Wen et al., 2024b), QED2 (Pailoor et al., 2023), and halo2-analyzer (Soureshjani et al., 2023) identify specific vulnerability patterns, while testing tools like MTZK (Xiao et al., 2025) and zkFuzz (Takahashi et al., 2025) use fuzzing to expose security issues.
-In this work, we focus on reducing the burden of writing ZKP programs by leveraging LLMs’ capabilities, complementing these efforts by helping developers audit and refine their code.
-7. Discussion, Limitation and Future Work
-ZKP Program Complexity.
-Program complexity for ZKPs is measured by the size of arithmetic circuits.
-It is a critical factor in the efficiency and scalability of ZKP systems, directly determining prover cost and overall performance.
-To reduce that, a range of compilers have been proposed to optimize circuit construction (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022), while new protocols reduce complexity through advanced arithmetizations with custom gates and lookup arguments (Chen et al., 2023; Gabizon et al., 2019).
+The software engineering community has recently begun to explore tools and methodologies for ZKP development, with significant emphasis on security.
+Static analysis tools such as Circomspect (Dahlgren, 2022), ZKAP (Wen et al., 2024a), QED2 (Pailoor et al., 2023), CCC-Check (Kolozyan et al., 2025), and halo2-analyzer (Soureshjani et al., 2023) identify specific vulnerability patterns, while testing tools (Xiao et al., 2025; Takahashi et al., 2025; Chaliasos et al., 2025) use fuzzing to expose security issues.
+These tools are invaluable for post-hoc analysis and verification, and can be integrated into our system for further security assurance.
+Our work complements these efforts by focusing on the generation phase, leveraging LLMs’ capabilities to reduce engineering burden.
+7. Discussion and Limitations
+ZKP Program Efficiency.
+Program efficiency for ZKPs is measured by the size of arithmetic circuits, which directly determines prover cost and overall performance.
+To reduce circuit size, a range of compilers have been proposed to optimize circuit construction (Stronati et al., 2024; Aztec Labs, 2025; Ozdemir et al., 2022), while new protocols reduce cost through advanced arithmetizations with custom gates and lookup arguments (Chen et al., 2023; Gabizon et al., 2019).
 On the programmer side, several techniques further improve efficiency, such as adopting ZK-friendly hash functions (Grassi et al., 2021), gate packing, and aggregating bits into single field elements.
-For LLM-based code generation, however, producing efficient circuits is more difficult than generating correct ones.
-As future work, we plan to survey optimization techniques and guide LLMs in incorporating these strategies.
+For LLM-based code generation, however, producing efficient circuits is more challenging than generating correct ones.
+As future work, we plan to survey optimization techniques and guide LLMs on this.
 LLM Training and Fine-tuning.
@@
 Our work takes a step toward mitigating this limitation by systematizing the generation of program examples in ZK DSLs, which can support targeted training and domain adaptation of LLMs.
-These examples may be used as a complementary resource for future research in ZK code generation.
-Threat to Validity.
-Our study on language knowledge relies on multiple-choice QA pairs that were initially generated with LLM assistance.
-Although some of these questions may contain leading hints or incorrect answers, we address this risk through repeated human validation and cross checking, ensuring that any residual noise has limited influence on the results.
-The evaluation in ZK-Eval employs carefully designed test cases that check both accepting and rejecting behaviors of generated programs and gadget usages.
-While test cases cannot fully guarantee correctness, they are constructed to include representative corner cases and to enforce soundness and completeness requirements, thereby reducing the likelihood of false positives.
-Furthermore, the gadget suite used by ZK-Eval and ZK-Coder consists of 35 representative components selected for their prevalence in ZK libraries and for their coverage of core patterns.
-This set strikes a balance between complexity and tractability, spanning arithmetic, boolean, comparison, etc., and has already demonstrated clear performance improvements in evaluation.
-Looking forward, the suite is designed for incremental extension, allowing more advanced and cryptographic composites to be incorporated in future work.
+These examples can serve as a complementary resource for future research in ZK code generation.
+Real-World Representativeness.
+We examine the representativeness and scope of our evaluation along two complementary dimensions.
+First, from a structural standpoint, the programs generated in § 5.4 exhibit an average cyclomatic complexity of 22.81, which exceeds the average observed in production-grade open-source ZK repositories (16.8 for Noir and 17.5 for Circom).
+This indicates that our benchmarks not only mirror but often surpass the structural depth of real-world ZK logic.
+Second, while our tasks are designed to isolate the core challenge of constraint specification, our evaluation is not “API-blind.” We recognize that LLM capabilities in standard and cryptographic API usage are already well established in prior work (Qin et al., 2023; Zhuo et al., 2024; Wang et al., 2025b; Zhu et al., 2025).
+Consequently, our test suite focuses on verifying that these APIs are integrated correctly within the circuit logic, covering a considerable number of real-world in-circuit invocations such as cryptographic primitives and data-processing utilities.
+This dual focus ensures that ZK-Coder is evaluated against the rigorous logical demands of production-level ZK development.
+Threats to Validity.
+We identify these potential threats.
+➀ Our language-knowledge study uses multiple-choice QA items that were initially generated with LLM assistance.
+To mitigate bias and errors, we performed iterative human validation and cross-checking; while some residual noise may remain, we expect it to have limited impact on the reported trends.
+➁ The ZK-Eval assessment uses hand-crafted test cases that exercise both accepting and rejecting behaviors of generated programs and primitive usages.
+Although tests cannot prove correctness, they include representative corner cases and explicit checks for soundness and completeness, which substantially reduce the risk of false positives.
+The high mutation detection rate reported in § 3.1.3, achieved through oracle-guided mutation testing, further supports this claim by demonstrating strong sensitivity to subtle constraint specification errors.
+➂ The reliability of our automated evaluation depends on test-case coverage.
+We mitigate this through extensive cross-validation and edge-case injection, achieving a high mutation score.
+While this is not a formal proof of logical equivalence, it provides strong empirical assurance of correctness.
+➃ The choice of prompting strategy may influence LLM performance.
+We employ standard zero-shot and few-shot prompting techniques without extensive prompt engineering or fine-tuning.
+To ensure experimental integrity, we verified that no test cases or ground-truth solutions were included in the prompts, preventing information leakage that could artificially inflate success rates.
+➄ Security is critical for ZKPs.
+While integration with auditing tools (Pailoor et al., 2023; Kolozyan et al., 2025) could further enhance assurance, code generation from natural language specifications may face inherent ambiguity and underspecification.
+Users must exercise caution and validate generated code before deployment in security-critical application.
 8. Conclusion
-We presented ZK-Eval, the first benchmark for systematically assessing LLM capability in ZK program synthesis, and showed that while models master language knowledge, they struggle with gadget construction and end-to-end synthesis.
-To address these gaps, we introduced ZK-Coder, a retrieval-augmented framework that guides LLMs from abstract constraint sketches to executable ZK code through constraint-guided retrieval and syntax-semantic repair.
-We show that ZK-Coder significantly improves the code generation capability of LLMs on both Circom and Noir.
-Appendix A Illustrative Workflow Example
-We demonstrate the workflow using ZK-Coder to generate a ZKP program that can be used to prove the correctness of a Sudoku assignment in Fig.
-13.
-Given a $4\times 4$ Sudoku grid, the goal is to verify that each row, column, and $2\times 2$ block contains all digits 1144 exactly once.
-Figure 13. ZK-Coder’s Illustrative Example Workflow on Proving Sudoku Correctness.
-In this illustrative workflow example, ZK-Coder first reads the problem description and formulates it in ZKSL, encoding the necessary constraints on the Sudoku grid.
-The checker validates the sketch and analyzes it to extract the set of constraint gadgets required.
-For each gadget, relevant implementation hints are retrieved, and an initial Circom implementation is generated.
-The compiler then checks this program for syntactic validity, returning diagnostics that guide interactive refinement.
-Once the code compiles, the model is prompted to generate a small test instance.
-If the resulting prover behavior deviates from expectations, the model is asked to repair the program.
-This repair loop continues until either the program passes the LLM-supplied tests or the retry budget is exhausted.
-We also present two concrete program examples generated by ZK-Coder for this Sudoku problem, one in Circom (Fig.
-14) and one in Noir (Fig.
-15).
-Both programs implement the same verification logic.
-We annotated the source code with comments to point out key sections, such as input signal declarations, gadget imports or implementations, and composing gadgets with field arithmetic operations to implement the constraints outlined in the ZKSL sketch.
-Figure 14. Example Circom Program Generated by ZK-Coder.
-Figure 15. Example Noir Program Generated by ZK-Coder.
-Appendix B Filtered Tasks in Dataset Curation
-As introduced in the main paper, our dataset construction is based on adapting existing programming tasks from HumanEval (Chen et al., 2021) and LiveCodeBench (Jain et al., 2024).
-To ensure that tasks are suitable for ZK DSLs (no floating-point numbers, no string manipulation, no I/O, deterministic output length, etc.), we filter the tasks using a combination of LLM assisted filtering and manual inspection.
-We now present the complete set of filtered tasks in Table 6 and Table 7.
-Table 6. Filtered tasks from HumanEval.
-Filtered Task IDs from HumanEval (Total 68 Tasks)
-3 5 8 9 13 24 31 33 35 36 37 39 40 41 42 43 46
-49 52 53 55 57 59 60 62 63 65 69 70 72 73 75 76 77
-83 85 88 90 94 97 102 106 108 109 110 114 115 116 121 122 123
-126 127 128 131 135 138 139 142 145 146 147 150 151 152 155 157 159
-Table 7. Filtered tasks from LiveCodeBench.
-Filtered Task IDs from LiveCodeBench (Total 34 Tasks)
-43 44 62 66 78 94 119 124 129 153 156 172 203 225 232 240 243
-250 261 265 280 290 297 305 322 323 336 346 353 367 373 375 384 389
-Appendix C Supplementary Results
-In this section, we present supplementary results that complement the main findings of the paper but were omitted earlier due to space limitations.
-C.1. Error Distribution For Gadget Implementation
-Previously, we reported the error distribution averaged across four models.
-Here, we provide the per-model error distributions.
-The breakdown by individual model is consistent with the overall trend: OpenAI’s reasoning models outperform open-source models, and the most common error types remain consistent across models.
-Figure 16. Error Distribution on GPT-o4-mini.
-Figure 17. Error Distribution on GPT-o3.
-Figure 18. Error Distribution on DeepSeek-V3.
-Figure 19. Error Distribution on Qwen3.
-Appendix D Complete Set of Constraint Gadgets
-The constraint gadgets used in the main paper during both measurement and augmentation.
-In measurement, they serve as the targets for evaluating LLMs’ ability to correctly employ or implement gadgets when expressing constraints.
-In augmentation, they are treated as primitive operators in the sketch language, and in constraint-guided retrieval we prepared knowledge snippets for each gadget to assist LLMs in code generation.
-The collection of gadgets is consolidated from three sources: (i) the official Circom standard library (iden3, 2025), (ii) the ZK-kit library (Nico Serrano, [n.
-d.]), and (iii) the Z3Py API (Research, [n.
-d.]).
-The complete list is shown in Table 8.
-We classify gadgets into five categories: Logical, Bitwise, Comparison, Arithmetic, and Composite, according to their functionality.
-The typing column specifies input and output types, which are used to disambiguate operators during constraint-guided retrieval.
-While this set is not exhaustive, it spans a wide range of common constraints in ZK programming.
-We expect to extend it with additional gadgets in future work.
-Table 8. List of constraint gadgets used in measurement and augmentation.
-Name Category Typing Source
-And Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
-And Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
-Or Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
-Or Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib,Z3Py
-XOr Logical Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
-XOr Logical Bool $\times\ldots\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
-Not Logical Bool $\rightarrow$ Bool Circomlib, Z3Py
-Or Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
-Or⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
-And Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
-And⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
-XOr Bitwise Field $\times$ Field $\rightarrow$ Field Z3Py
-XOr⋄∗ Bitwise Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
-Equal Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
-NotEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
-LessThan Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
-LessThanOrEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
-GreaterThan Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
-GreaterThanOrEqual Comparison Field $\times$ Field $\rightarrow$ Bool Circomlib, ZK-Kit, Z3Py
-Add Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
-Subtract Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
-Multiply Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
-Divide Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
-Modulo Arithmetic Field $\times$ Field $\rightarrow$ Field ZK-Kit, Z3Py
-Power Arithmetic Field $\times$ Field $\rightarrow$ Field Z3Py
-FloorDivide Arithmetic Field $\times$ Field $\rightarrow$ Field ZK-Kit, Z3Py
-Negate Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
-Absolute Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
-Inverse Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
-Sign Arithmetic Field $\rightarrow$ Field Circomlib, Z3Py
-Conditional Composite Bool $\times$ Field $\times$ Field $\rightarrow$ Field Circomlib, Z3Py
-Conditional Composite Bool $\times$ Bool $\times$ Bool $\rightarrow$ Bool Circomlib, Z3Py
-Sum Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
-Product Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
-Distinct Composite Field $\times\ldots\times$ Field $\rightarrow$ Field Z3Py
+We presented ZK-Eval, the first benchmark for assessing LLM capability in ZK program generation, revealing that models struggle with composing algebraic primitives for verification constraints.
+To address this, we introduce ZK-Coder, which guides LLMs from constraint sketches to executable code through sketch-guided retrieval and iterative refinement, significantly improving code generation for both Circom and Noir.